### v2

# Almost commuting self-adjoint operators and measurements

Abstract
We study the problem when an -tuple of self-adjoint operators in an infinite dimensional separable Hilbert space with small commutators is close to an -tuple of commuting self-adjoint operators on We give an affirmative answer to the problem when the synthetic-spectrum and the essential synthetic-spectrum are close. In the case that we show that a pair of almost commuting self-adjoint operators in an infinite dimensional separable Hilbert space is close to a commuting pair of self-adjoint operators if and only if a corresponding Fredholm index vanishes outside of an essential synthetic-spectrum. Examples are also exhibited that, in general, the answer to the problem when is negative even the associated Fredholm index vanishes. This is an attempt to solve a problem proposed by David Mumford related to quantum theory and measurements.
1 Introduction
In the First International Congress of Basic Science held in 2023, David Mumford gave the opening plenary lecture on Consciousness, robots and DNA ([32]). At the very beginning, he cautioned the audience that there would be no mathematics in his talk. Nevertheless, by the very end of this magnificent lecture, when he ended his talk on DNA being a measuring instrument opening a Pandoraβs box, he proposed a problem in -algebras about non-commuting self-adjoint operators to be approximated by commuting ones.
A version of this problem proposed by Mumford may be reformulated below: Let and be a positive integer. Is there a constant such that the following statement holds? If is any separable Hilbert space and are self-adjoint operators on (with ) such that
| (eβ1.1) |
there exist self-adjoint operators on such that
| (eβ1.2) |
In the case that and is any finite dimensional Hilbert space (no bound on the dimension) the exactly the same problem is known as von-Neumann-Kadison-Halmos problem for almost commuting self-adjoint matrices (see [11] and [12]). In that case, the problem has an affirmative solution [20] (see also [8] and [13], and, for a fixed dimension, see [31]). In infinite dimensional Hilbert space, as observed by Mumford in his new book (p.182), the answer in general is negative (see also Example 4.6 of [8]). The affirmative solution for almost commuting self-adjoint matrices led the further study in the subject of weakly semi-projectivity in -algebras. It also has a deep impact on the Elliott program of classification of amenable -algebras (see [7], [24], [26] and [28]). The Mumfordβs problem opens a new door to the future applications of this part of -algebraβ theory in previous unexpected areas such as DNA and artificial intelligence ([33]).
The first main result of this paper is as follows:
Theorem 1.1.
Let and There exists satisfying the following: Suppose that is an infinite dimensional separable Hilbert space and are bounded self-adjoint linear operators on with () such that
| (eβ1.3) |
where and Then there are bounded self-adjoint linear operators on such that
| (eβ1.4) |
(See Remark 6.5 for the case that each is compact.)
Here and are the sets of -synthetic-spectrum, and essential -synthetic-spectrum, respectively (see Definition 2.7). In (eβ1.3), is the Hausdorff distance between the two compact subsets of the -dimensional unit cube. In the special case that is full, i.e., it contains the whole then the second part of (eβ1.3) is automatically satisfied.
Remark 6.6 provides some justification of the second condition related to macroscopic observables and measurements, and a discussion of the Mumford problem.
More generally, let be a Hilbert (right) module over a -algebraβ be the -algebraβ of all bounded module maps on with adjoints and let be the linear span of rank one module maps of the form (for all ), where Denote by the closure of which is a -algebra. We also have the following:
Theorem 1.2.
Let and There exists satisfying the following: Suppose that is a countably generated Hilbert-module over a -unital purely infinite simple -algebraβ and are self-adjoint bounded module maps with () such that
| (eβ1.5) |
Suppose also
(i) if or
(ii) if and
where and
and is the quotient map.
Then there are such that
| (eβ1.6) |
For a pair of self-adjoint operators, we also have a more definite answer for some Hilbert modules.
Theorem 1.3.
Let There exists satisfying the following:
Suppose that
(i) is an infinite dimensional separable Hilbert space, or
(ii) is a countably generated Hilbert module over a -unital purely infinite simple -algebraβ
such that has real rank zero, or
(iii) is a countably generated Hilbert module over a -unital simple
-algebraβ of stable rank one such that has real rank zero and is simple.
Suppose also that are self-adjoint bounded module maps on with () such that
| (eβ1.7) |
for all where is the quotient map. Then there are self-adjoint bounded module maps such that
| (eβ1.8) |
Moreover, the condition that has real rank zero can be removed if we assume
that
is connected.
The second condition in (eβ1.7) means that a Fredholm index vanishes (see Definition 2.4). Note that both case (ii) and case (iii) are generalization of case (i). In case (ii), is stable (whenever is not a finitely generated projective Hilbert module), and is simple (like the Calkin algebra). In case (ii), we keep the feature that is simple and has real rank zero.
Case (i) in Theorem 1.3 is a more definite result than that of Theorem 1.1 in the case We actually prove a slightly more general result (see Theorem 5.3 below). One should also compare (i) in Theorem 1.3 to Theorem 1.1 of [17] (see Remark 5.6). In the case that however, the normal elements may not be approximated by normal elements in with finite spectrum. We also show that condition of the vanishing index associated with the pair of operators is necessary. Since it is well-known that spectrum (and the essential spectrum) of a bounded operator is notoriously unstable regarding perturbation, it is necessary to consider some version of -spectrum (which is usually larger than the spectrum). For we introduce the notion of -synthetic-spectrum for -tuples of self-adjoint elements in a unital -algebraβ (see Definition 2.7). We will also show that, unfortunately, the version of Mumfordβs problem formulated above has a negative solution in general even the associated Fredholm index vanishes.
In quantum mechanics, one often heard βwhen the commutators go to zero, we recover the classical systemβ. If βrecoverβ means the observables are approximated by commuting observables, then example in (iii) of Chapter 14 of [33] (Example 4.6 of [8]), Proposition 5.5 and Proposition 6.4 all show that, mathematically, there are serious (topological) obstacles prevent this type of statements from being true. Nevertheless Theorem 1.1 with some interpretation in Remark 6.6 may provide an alternative statement.
The paper is organized as follows. Section 2 contains some notations used in the paper and gives the definition of -synthetic-spectrum and -near-spectrum for -tuples of self-adjoint operators, and discusses the relationship between them (see also Proposition 2.16). Section 3 is a study of -tuples of almost commuting self-adjoint elements in a unital purely infinite simple -algebra. Section 4 presents the proof of Theorem 1.1 and Theorem 1.2. In Section 5, we will discuss the case of a pair of almost commuting self-adjoint bounded module maps and provide a proof of Theorem 1.3. In section 6, the last section, we show why the Fredholm index does not appear in the statement of Theorem 1.1. We then present an example that an -tuple version of Theorem 1.3 does not hold in general. A final remark was added which contains the justification of the second condition in (eβ1.3) and its connection to the Mumford original question.
2 Preliminaries
Definition 2.1.
Let be a -algebra. Denote by the set of all self-adjoint elements in Denote by the multiplier algebra (the idealizer of in ). Denote by and the group of invertible elements and the path connected component of containing the identity of respectively.
If and we write
| (eβ2.9) |
Definition 2.2.
Let be an infinite dimensional separable Hilbert space. Denote by the -algebraβ of all bounded linear operators on and by the -algebraβ of all compact operators on Denote by the quotient map.
Let Recall that the essential spectrum of is the spectrum of
Definition 2.3.
Let be a -algebraβ and be a Hilbert (right) module over (see [15]), or Hilbert -module. Denote by the -algebraβ of all bounded module maps on with adjoints. A rank one module map is a bounded module map of the form for all (and fixed ), where is the -value inner product on Denote by the linear span of rank one module maps and the (norm) closure of Following Theorem 1 of [15], we identify with
Definition 2.4.
Let be a unital -algebraβ and be an invertible element. Define Note that is a unitary. Define
| (eβ2.10) |
Then if and only if Let be such that is invertible in Then
| (eβ2.11) |
Definition 2.5.
In by we mean the euclidian distance between and
Denote by If and define
Let for all be the constant function,
be defined by
for
This notation will be used throughout this paper.
Note that is generated by
Definition 2.6.
Fix an integer Define
has only finitely many points.
Definition 2.7.
Let In what follows, for each we choose a fixed integer such that Denote We write Then is -dense in Moreover, if
Let be a unital -algebraβ and be an -tuple of self-adjoint elements in with (). Fix Let be such that if if and is linear in and in Define, for
| (eβ2.12) | |||||
| (eβ2.13) |
Note that we do not assume that mutually commute and the product in (eβ2.13) has a fixed order.
For write We may write and Set
| (eβ2.14) |
This union of finitely many closed balls is called -synthetic-spectrum of the -tuple which, by the definition, is compact.
Let Suppose Then Note that (). Hence It follow that, if then Then, if
| (eβ2.15) |
Note also that, if then
| (eβ2.16) |
To see this, let There is with such that There is such that Write and If then Therefore Hence Therefore
If for some infinite dimensional separable Hilbert space, define
| (eβ2.17) |
which will be called essential -synthetic-spectrum. It is clear that
| (eβ2.18) |
Definition 2.8.
Let and be -algebras. A linear map is a c.p.c. map if it is completely positive and contractive.
Definition 2.9.
Let be a unital -algebraβ and for some Let us assume that Fix Suppose that is a compact subset and is a unital c.p.c. map such that
| (eβ2.19) | |||||
| (eβ2.20) | |||||
| (eβ2.21) |
for any which has value on an open ball with the center (for some ) and the radius Then we say that is an -near-spectrum of the -tuple We write
If, moreover, is a unital homomorphism, then we say is an - spectrum of the -tuple
Let be an -near-spectrum and be a -near -spectrum for respectively. Suppose that then, by the definition, is also an -near-spectrum. In particular, if then
Let be an infinite dimensional separable Hilbert space and let If is an - spectrum of we say that is an essential - spectrum for In this case we write
Remark 2.10.
We often use the -near spectrum as it associated with a c.p.c. map, whenever the -tuple sufficiently almost commutative as shown in the next proposition (2.11). But it is not unique and depends on the choice of the c.p.c. map. On the other hand, the -synthetic-spectrum is always defined and in the computation world, it is computable. Their relationship is described in Proposition 2.15 below (see also Proposition 2.16). In practice, it is quite possible that it is more difficult to compute the spectrum of an operator where than to compute the norm of
Proposition 2.11.
Fix For any there exists satisfying the following: Suppose that is a unital -algebraβ and with such that
| (eβ2.22) |
Then
(1) and
Proof.
Suppose that, for case (1), (or for case (2), ). Then we obtain an and a sequence of unital -algebras and a sequence of -tuples with (), such that
| (eβ2.23) |
and, (or, for case (2), )
Consider and quotient map Let Then Let be the -subalgebraβ of generated by Then for some non-empty compact subset of Let be the monomorphism given by with
To show (1), consider Fix Choose such that then Hence, since is a monomorphism,
| (eβ2.24) |
Recall that Hence there exists such that
| (eβ2.25) |
It follows that for all In other words, has -synthetic-spectrum containing This proves (1).
For case (2), we obtain, by the Choi-Effros Theorem ([4]), a c.p.c. map such that Write where each is a c.p.c map, Then
| (eβ2.26) |
Let be an -dense subset, and with if and if
Lemma 2.12.
Let be a compact metric space and be a finite generating subset.
Then, for any and any finite subset there exists satisfying the following: Suppose that is a unital -algebraβ and are unital c.p.c. maps such that
| (eβ2.29) |
Then, for all
| (eβ2.30) |
Moreover, if and then, for any and and finite subset we may also require that, for all
| (eβ2.31) |
Proof.
Suppose the lemma is false. Then there exist a finite subset and a sequence of unital -algebras and two sequences of unital c.p.c. maps such that
| (eβ2.32) | |||
| (eβ2.33) | |||
| (eβ2.34) |
Moreover, in the case that and there exists and a finite subset such that
| (eβ2.35) |
for all
Put Let the quotient map. Then, since is a generating set of by (eβ2.33), is a unital homomorphism, and by (eβ2.32), Hence
| (eβ2.36) |
But this implies that, for all sufficiently large
| (eβ2.37) |
This leads a contradiction to (eβ2.34). The first part of the lemma then follows.
For the second part, since is a unital homomorphism, for any
| (eβ2.38) |
It follows that, for all large and
| (eβ2.39) |
Another contradiction. So the lemma follows. β
Corollary 2.13.
Fix and There exists satisfying the following: Suppose that is a unital -algebra, with (), is a compact subset, and is a unital c.p.c. map such that
| (eβ2.40) |
Then, for
| (eβ2.41) |
Proof.
Let be as defined in 2.7, where (). There is satisfying the following: for any unital -algebraβ and any pair of self-adjoint elements with if then
| (eβ2.42) |
where is defined in 2.7.
Let and Choose such that (the second part of) Lemma 2.12 holds for above and (in place of ). Choose such that, for any unital c.p.c. map (any unital -algebraβ ), that
| (eβ2.43) |
implies, for
| (eβ2.44) |
Choose
Suppose that and satisfy the assumption of this lemma for Since and
| (eβ2.45) |
for any -tuple of self-adjoint elements in with (), by the choice of (and ) and by the second part of (eβ2.40) and (eβ2.42), we have, for each
| (eβ2.46) |
By the choice of we also have
| (eβ2.47) |
By the choice and applying (the second part of) Lemma 2.12, we have
| (eβ2.48) |
It follows that (keeping in mind of (eβ2.45))
| (eβ2.49) |
Combining (eβ2.49) with (eβ2.47), we obtain
| (eβ2.50) |
We conclude that, by (eβ2.50) and (eβ2.46),
| (eβ2.51) | |||
| (eβ2.52) | |||
| (eβ2.53) |
β
Definition 2.14.
Let be a metric space (we only consider the case that ). Recall that the Hausdorff distance of two compact subsets of is defined by
| (eβ2.54) |
Let be the set of all non-empty compact subsets of Then is a compact metric space with the metric
Proposition 2.15.
Fix For any there exits satisfying the following:
(1) Suppose that is a unital -algebraβ and with () such that has a non-empty -near-spectrum If is also a non-empty -near-spectrum of and then
| (eβ2.55) |
(2) Suppose, in addition to (1), has a (closed two-sided) ideal is the quotient map, and and Then
| (eβ2.56) |
Proof.
We may assume that We first prove (1).
Let be an -dense subset of Choose with if and if Put
Note that is a generating set of Let be given by Lemma 2.12 (for and ) and be given by Corollary 2.13. Choose
Now suppose that and are -near-spectra of the -tuple Let be c.p.c maps such that
| (eβ2.57) | |||
| (eβ2.58) | |||
| (eβ2.59) |
for any with for all whenever (if ) or (if ). Let We may assume that Then By (eβ2.59),
| (eβ2.60) |
By (eβ2.58) and the choice of applying Lemma 2.12,
| (eβ2.61) |
Since if we have Therefore
| (eβ2.62) |
This holds for all Exchanging the role of and we obtain
| (eβ2.63) |
Let and
By the choice of and applying Corollary 2.13, we have that
| (eβ2.64) |
Pick There is such that Recall that if As we have Hence for all It follows that
| (eβ2.65) |
Thus, by (eβ2.64),
| (eβ2.66) |
This implies that
| (eβ2.67) |
Next, let By (eβ2.64),
| (eβ2.68) |
Hence Thus
| (eβ2.69) |
It follows that
| (eβ2.70) |
Hence
| (eβ2.71) |
For (2), let us keep notation above. Then is immediate, and that and that follow from part (1). It remains to show that
There is a unital c.p.c. map such that
| (eβ2.72) | |||
| (eβ2.73) | |||
| (eβ2.74) |
for any with if for some Choose We may assume that Hence It follows from (eβ2.74) that
| (eβ2.75) |
On the other hand, by the choice of (eβ2.58), (eβ2.73), and applying Lemma 2.12, we have
| (eβ2.76) |
Thus
| (eβ2.77) |
Hence It follows that
| (eβ2.78) |
Hence β
Proposition 2.16.
Let There exists satisfying the following: Suppose that is a unital -algebraβ and such that (), and
| (eβ2.79) |
Then
Proof.
Otherwise, one obtains a sequence of unital -algebras a sequence of elements with such that
| (eβ2.80) |
and a sequence of numbers Put and Note that and Since is compact, without loss of generality, we may assume that for some compact subset in as and
Let and be the quotient map. Put Then, by (eβ2.80), is normal. Let Then defined by for is a unital monomorphism. By the Choi-Effros Lifting Theorem ([4]), there is a sequence of unital c.p.c. maps such that Let be as in Proposition 2.15. Then, since is a homomorphism, for all large is a -near-spectrum for Applying Proposition 2.15, we may assume that, for all
| (eβ2.83) |
Note that, by (eβ2.82) and (eβ2.83), Let Note that Hence there are such that ()
| (eβ2.84) |
Hence, for all large
| (eβ2.85) |
This is a contradiction. So the lemma follows. β
3 Purely infinite simple -algebras
Recall that the Calkin algebra (when is infinite dimensional separable Hilbert space) is purely infinite and simple. The results in the section will be used in the next four sections.
Definition 3.1.
An even - brick with corner is a subset of of the form
| (eβ3.86) |
For the remaining of this paper, all -bricks are even and has the corner An open -brick is the interior of the -brick By boundary of the -brick we mean
A -brick combination is a finite union of -bricks. The -brick combination has the following easy property: If are two distinct -bricks of then
We think Proposition 3.2 are well known.
Proposition 3.2.
If is a union of finitely many compact convex subsets in (for some integer ), then is finitely generated,
Proof.
We prove this by induction on the number of convex sets.
If is convex, is contractive. Hence and
Suppose that the proposition is proved for many closed convex sets,
Now suppose that is a union of many compact convex sets in Write where each is a compact convex subset of
Put Then which is a union of many compact convex subsets. By the induction assumption, is finitely generated ().
Then
| (eβ3.87) |
is a pull back. By the Mayer-Vietoris sequence in -theory for -algebras (see [14]), we obtain the following commutative diagram.
| (eβ3.91) |
Since we have established that and are all finitely generated, we conclude that is also finitely generated. This ends the induction.
β
Definition 3.3.
Let be a compact subset of Define to be the union of all even -bricks (in with corners at ) which has a non-empty intersection with
We will use the following easy facts:
(i) (ii) If is one of the bricks in then (iii) For any and (iv) is finitely generated (this follows from Proposition 3.2).
In what follows is called the -brick cover of (The introduction of even -bricks is for convenience not essential.)
Definition 3.4.
Let be a compact metric space. We say that has property (F), if is a disjoint union of finitely many path connected compact metric spaces such that is finitely generated (), Choose a base point Write Denote by
Let be a -algebraβ and a homomorphism. Denote by the restriction.
Every finite CW complex has property (F). If is a -brick combination in then (by Proposition 3.2) also has property (F).
Recall that if has finite dimensional range, then there are mutually orthogonal projections and such that
| (eβ3.92) |
Theorem 3.5.
Let be a compact metric space with property (F). Suppose that is a unital purely infinite simple -algebraβ and is a unital injective homomorphismβ such that
| (eβ3.93) |
Then, for any and any finite subset there exists a set of finitely many mutually orthogonal projections with and such that
| (eβ3.94) |
where for all Moreover, if is connected, we may require that either or in
Proof.
Since has property (F), we may write where each is path connected. Let be such that and if Put (). By considering for each we reduce the general case to the case that has only one path connected component. Put for some Then we have a splitting short exact sequence
Recall that is finitely generated (). It follows that, for any unital -algebraβ
| (eβ3.95) |
Define by for all Then one computes that, by (eβ3.93),
| (eβ3.96) |
Fix and finite subset Applying Theorem A of [6], we obtain an integer a unitary and such that, for all
| (eβ3.97) |
Define by for all Then we may rewrite for all where is a set of mutually orthogonal projections such that Choose mutually orthogonal projections such that where and in (recall that is purely infinite simple). Define by for all By replacing by and by we may assume, without loss of generality, that and
| (eβ3.98) |
(But then is a unitary in ).
Recall that, by S. Zhangβs theorem ([38]), has real rank zero. Since is injective, by applying Lemma 4.1 of [7], we obtain a set of mutually orthogonal non-zero projections with such that
| (eβ3.99) |
There are such that in (recall again that is purely infinite simple). We may then write
| (eβ3.100) |
where for all There are such that
| (eβ3.101) |
Let Then and Moreover,
| (eβ3.102) |
Hence
| (eβ3.103) |
Define for all Then is a unital homomorphismβ from to with finite dimensional range and, by (eβ3.103), (eβ3.100) and (eβ3.98), for all
Finally, since in as we arranged earlier, by the definition of and the βMoreoverβ part of the statement also holds. β
Let be a compact subset. We use the usual Euclidian metric. Recall that, for any
| (eβ3.104) |
Theorem 3.6.
Fix For any and finite subset There exists satisfying the following:
Let be a compact subset of Suppose that is a unital purely infinite simple -algebraβ and is a unital injective homomorphism. If is any compact subset with property (F) such that where then there is a a unital injective homomorphismβ such that
| (eβ3.105) |
where is the quotient map by restriction ( for all ).
Proof.
Let and finite subset be given. Choose such that, for all
| (eβ3.106) |
Fix any Suppose that is a compact subset of such that Since is compact, there is with such that
| (eβ3.107) |
Put and Then There is a finite -net such that where Then
Choose a non-zero projection with Since is a purely infinite simple -algebra, there is a unital embedding Applying Theorem 2.8 of [16], one obtains a unital embedding Note in There is a partial isometry such that and Define by for all Then is a unital injective homomorphism.
Choose mutually orthogonal non-zero projections such that Define by Define by Then is injective. One computes that
| (eβ3.108) |
It follows from Theorem 1.7 of [6] that there is a unitary such that
| (eβ3.109) |
Applying Theorem 3.5, there is a unital homomorphismβ with finite dimensional range such that
| (eβ3.110) |
We may write for all where and are mutually orthogonal non-zero projections with By choosing a better approximation if necessary, without loss of generality, we may assume that is -dense in Recall that is -dense in and -dense in For each choose exactly one such that Then, we may assume where is a set of distinct points such that
| (eβ3.111) |
Note that we view Then set Put where for all
By the choice of we have, for all
| (eβ3.112) |
Recall that Hence in There are partial isometries such that and Define Then
| (eβ3.113) | |||||
| (eβ3.114) |
So is a unitary in Note that
| (eβ3.115) |
Then, combining with (eβ3.110) and (eβ3.112), for all
| (eβ3.116) |
Define by
| (eβ3.117) |
Then is a unital injective homomorphism. Moreover, for all by (eβ3.109), the definition of above and by (eβ3.116),
| (eβ3.118) | |||||
To see the second formula in (eβ3.105), we may write, from (eβ3.117), since factors through
| (eβ3.119) |
β
Corollary 3.7.
Fix an integer and
Suppose that is a compact subset of and suppose that is a unital purely infinite simple -algebraβ and is a unital injective homomorphism. If is any compact subset with property (F) such that where then there is a unital injective homomorphismβ such that
| (eβ3.120) |
where is the quotient map by restriction ( for all ).
Proof.
In the proof of Theorem 3.6, we may choose to be a -brick combination with (the same in the proof of Theorem 3.6). Thus we may state a variation of Theorem 3.6 as follows.
Corollary 3.8.
Let be a compact metric space which is a a compact subset of (for integer ) and be a unital purely infinite simple -algebra. Suppose that is a unital injective homomorphism. Then, for any and any finite subset there exists such that, for any integer with there exists a unital injective homomorphismβ such that
| (eβ3.122) |
where is the quotient map by restriction, is a finite subset such that (and where is the -brick cover of ).
Theorem 3.9.
Let and There is satisfying the following: If is a unital purely infinite simple -algebraβ and with () such that
| (eβ3.123) |
then there is integer with and and a -brick combination and a unital injective homomorphismβ such that
| (eβ3.124) |
Proof.
Let It follows from Corollary 1.20 of [22] that there exists such that when there are such that
| (eβ3.125) |
Let be the commutative -subalgebraβ generated by Then there exists a unital homomorphismβ such that Hence for some compact subset of Denote the isomorphism give above. We may assume that
Put Choose any with and By Theorem 3.6, we obtain a -brick combination and a unital injective homomorphismβ such that
| (eβ3.126) |
Hence
| (eβ3.127) |
β
4 Multiple self-adjoint operators
The main purpose of this section is to prove Theorem 4.6.
The following lemma is a known (see the proof of Theorem 3.15 of [3]).
Lemma 4.1.
Let be a unital -algebraβ and be an ideal of which is a -unital -algebraβ of real rank zero. Suppose that is a projection such that where is the quotient map. Then there is a projection such that
Lemma 4.2.
Fix For any there exists satisfying the following: Suppose that is a unital -algebra, is a separable ideal of real rank zero, is the quotient map such that is purely infinite simple, and with () such that the -tuple has a -near-spectrum and has -spectrum Then there are mutually orthogonal non-zero projections a -dense subset in and a c.p.c. map where such that
| (eβ4.128) | |||
| (eβ4.129) |
and
Proof.
Let be given. and Choose Since is a -near-spectrum of there is a unital c.p.c. map such that
| (eβ4.130) |
Since is a -spectrum of there exists a unital monomorphism such that
| (eβ4.131) |
Let be a non-zero projection such that Choose a partial isometry such that
| (eβ4.132) |
It follows that in Let be a -dense subset of Let be mutually orthogonal non-zero projections in such that in Define by
| (eβ4.133) |
Note that in Then, since is purely infinite and simple, by Theorem 1.7 of [6], there exists a unitary such that
| (eβ4.134) |
Let and
Then we may write
| (eβ4.135) |
Since in by applying Lemma 4.1, we obtain projections such that
| (eβ4.136) | |||
| (eβ4.137) |
Let Then is a projection. By the Choi-Effros lifting theorem ([4]), there is a c.p.c.map such that
Let be an approximate identity for and an approximate identity for consisting of projections. Put
Then is an approximate identity for and is an approximate identity for Choose such that
| (eβ4.139) |
Hence, since
| (eβ4.140) |
Put Then, by (eβ4.138), for
| (eβ4.142) | |||||
Similarly, for
| (eβ4.144) | |||||
Put Then
| (eβ4.145) | |||
| (eβ4.146) | |||
| (eβ4.147) |
Put for Then it is a c.p.c. map. Moreover
| (eβ4.149) | |||||
Hence, using the inequality above and (eβ4.130),
| (eβ4.150) |
To arrange we may split We write where are two mutually orthogonal projections and both are not in as Then define by for all We then replace with and replace with
The following proposition is a list of easy facts. We refer to Definition 2.1 of [10] for the definition of these properties. It is certainly known that has these properties and some of these may also stated in various places. We put here for our convenience.
Proposition 4.3.
Let be a -unital -algebra. Then has the following property
(1) --cancellation with for all
(2) --cancellation with for all
(3) -stable rank 1;
(4) -stable rank 1;
(5) stable exponential length
(6) stable exponential rank
(7) -divisible rank (for );
(8) -divisible rank (for ), and
(9) exponential length divisible rank (for all ).
Proof.
First, by [5], We also have, for any For any projection we have This immediately implies (1), (4) and (7).
It follows from Theorem 1.1 of [39] that and Therefore, (2), (3), (5), (6), (8) and (9) hold. β
Definition 4.4.
Let be a unital -algebra. Let and be maps, and We say satisfies condition (R, r, b, T, s) if has --cancellation with for all -stable rank -divisible rank and for all
If is a unital purely infinite simple -algebra, then satisfies condition for any and and If has stable rank one and real rank zero, then satisfies condition for any and
We will apply the following theorem which is a variation of that of Theorem 3.1 of [10].
Theorem 4.5.
Let where is a compact subset. Let be maps and
For any there exist a positive number and an integer satisfying the following: Suppose that is a unital -algebraβ satisfying condition such that for any compact metric space and are unital two c.p.c. maps such that
| (eβ4.151) |
then there are a unitary and a homomorphism with finite dimensional image such that
| (eβ4.152) |
Proof.
Let and Since we assume that for any compact metric space for any finite subset as in Theorem 3.1 of [10], we have that for any sufficiently almost multiplicative c.p.c. maps Then, by Proposition 4.3 and applying Theorem 3.1 of [10], we obtain and a finite subset such that, if
| (eβ4.153) |
there exists an integer a unital homomorphismβ and a unitary such that
| (eβ4.154) |
(Note as mentioned above, that whenever it makes sense, since for any compact metric space )
Theorem 4.6.
Let and Let and be maps and There exists satisfying the following: Suppose that is a unital -algebraβ such that for any compact metric space () satisfying condition and is an essential ideal which has real rank zero such that is purely infinite simple, for any projection and suppose that with () such that
| (eβ4.156) |
where and where is the quotient map. Then there are such that
| (eβ4.157) |
Proof.
Fix Let Fix a finite -dense subset such as for some large (as in 2.6). Assume that there are many points in There are at most many compact subsets of such that, for any compact subset there exists such that
| (eβ4.158) |
We will apply Theorem 4.5. Let (in place of ) be given by Theorem 4.5 for and (in place of ), Define
Put Let (in place of ) be given by Lemma 4.2 for (in place of ). Choose Put
Let be given by Proposition 2.15 for (in place of ). Choose Let be given by Proposition 2.11 for (in place of ) and let be given by Theorem 3.9 for
Let
Now suppose that with () such that
| (eβ4.159) |
By the choice of (and ), by applying Proposition 2.11, and are not empty set (see also (eβ2.15)) and the -tuple has an -near-spectrum and, by Theorem 3.9, has an -spectrum Moreover (by the choice of ) such that (by applying Proposition 2.15),
| (eβ4.160) |
By the choice of applying Lemma 4.2, we obtain mutually orthogonal nonzero projections a -dense subset a c.p.c. map such that (with )
| (eβ4.161) | |||
| (eβ4.162) |
and Put for all
Note that, since Choose and define by
By the choice of and (eβ4.162), as and applying Theorem 4.5, we obtain a unital homomorphismβ with finite dimensional range and a unitary such that
| (eβ4.163) |
Put
| (eβ4.164) |
Since has a finite dimensional range, there are mutually orthogonal projections (for some integer ) such that
| (eβ4.165) |
where
By the assumption that and (eβ4.160), is -dense in Note that
| (eβ4.166) |
Hence is -dense in Let
| (eβ4.167) | |||||
| (eβ4.168) |
By induction, since is -dense in we obtain such that and where
Put Define
| (eβ4.169) |
Then
| (eβ4.170) |
In we have a partial isometry such that
Put Then
| (eβ4.171) |
It follows that
| (eβ4.172) |
Put
| (eβ4.173) |
for all Then, by (eβ4.163), (eβ4.164), (eβ4.170), (eβ4.172) and (eβ4.161), for
| (eβ4.174) |
Put
| (eβ4.175) |
Then, since by (eβ4.174),
| (eβ4.176) |
Since is a homomorphismβ so is Hence
| (eβ4.177) |
β
Proof of Theorem 1.1
Proof.
4.7.
Proof of Theorem 1.2
Proof.
Let be a -unital purely infinite simple -algebraβ and be a countably generated Hilbert -module. Let Then, by Kasparovβs absorbing theorem (Theorem 2 of [15]), is an orthogonal summand of and hence is isomorphic to a hereditary -subalgebraβ of which is purely infinite simple. If is unital, by Theorem 1 of [15], Then the conclusion of Theorem 1.2 follows from Theorem 3.9 (without referring to synthetic spectra).
Otherwise, since is countably generated, is non-unital but -unital (see Proposition 3.2 of [2]). By Theorem 1 of [15], Note that (see Remark 2.5 (c) of [30]) which has real rank zero by [36]. It follows from Theorem 3.3 of [37] (see also Remark 2.5 (b) of [30]) that is purely infinite simple, and every projection in is equivalent to Thus, by Proposition 4.3, Theorem 4.6 applies. β
5 A pair of almost commuting self-adjoint operators on Hilbert modules
Definition 5.1 (Definition 3.1 of [8]).
For a unital -algebraβ denote by the set of elements with the property that for no ideal of is one-sided and not two-sided invertible in Recall that is said to have property (IR) if all elements in belongs to the norm closure of the group of invertible elements.
A non-unital -algebraβ has property (IR), if the -algebraβ obtained by adjoining a unit to has property (IR).
It follows from Lemma 4 of [23] that, for any projection has (IR).
Every -algebraβ of stable rank one has property (IR), and every purely infinite simple -algebraβ has property (IR).
Lemma 5.2.
Let There exists satisfying the following: Let be a unital purely infinite simple -algebra, be self-adjoint elements such that and
| (eβ5.178) |
Then there exists a pair of commuting self-adjoint elements such that
| (eβ5.179) |
whenever
(Note that the identity in (eβ5.179) also implies that )
Proof.
It follows from Theorem 3.9 that exists such that, whenever
| (eβ5.180) |
there exists a compact subset and a unital injective homomorphismβ such that
| (eβ5.181) |
Put and Then is a normal element in and Suppose that such that Then
| (eβ5.182) |
Hence
| (eβ5.183) |
It follows that Hence and are in the same path connected component of It follows that β
Theorem 5.3.
Let There exists satisfying the following: Suppose that is a unital -algebraβ with an essential ideal which is a -unital -algebraβ with real rank zero and with property (IR) such that is purely infinite simple. Suppose that with () such that
| (eβ5.184) |
for all where is the quotient map. Then, if either has real rank zero, or is path connected, there are such that
| (eβ5.185) |
Proof.
Fix Let be given by Theorem 4.4 of [8] for in place of
Choose Let be given by Proposition 2.15 for (in place of ) and Choose Let be given by Proposition 2.16 for (in place of ). Let be given by Theorem 5.2 for (in place of ).
Choose
Now assume that satisfy the assumption for Put
Recall that is a purely infinite simple -algebra. Therefore, by the choice of and applying Theorem 5.2 we obtain a compact subset a unital injective homomorphismβ such that
| (eβ5.186) |
for those for which where and In particular, has an On the other hand, by the choice of we have
| (eβ5.187) |
We claim: Suppose that is a bounded path connected component and that there is such that Then, for any
| (eβ5.188) |
To see the claim, we note that, by (eβ5.187), Then, by (eβ5.186) and by the assumption (eβ5.184),
| (eβ5.189) |
It follows that, for any
| (eβ5.190) |
and the claim follows.
Choose such that
Note that, if is path connected, then is also path connected.111 If for some then there is a path in connecting to and there is path connecting to in Let Choose the -brick cover of (see Definition 3.3). Note that if is path connected then is also connected.
Put Then
By Corollary 3.7, there is a unital injective homomorphismβ such that
| (eβ5.191) |
where is the quotient map by restriction. Put and Then
Denote by the unbounded path connected component of Let be a bounded path component. Then there are two cases:
(ii) if then
To see this, let us assume it is not the case (i). Then is in a bounded path connected open component ButοΌ since one must have Hence
Next we will show that To see this, let where is a bounded path connected component of Put
| (eβ5.192) |
If (i.e., it is case (i)), then, for
| (eβ5.193) |
Hence, by the second part of (eβ5.191),
| (eβ5.194) |
If and (i.e., in the case (ii)), by the second part of (eβ5.191) and the claim above, one computes that
| (eβ5.195) |
Hence
| (eβ5.196) |
Since and is finitely generated, is free. Therefore in Then, by Theorem 3.5, there is a normal element with which has finitely many points in the spectrum contained in such that
| (eβ5.197) |
Then, by (eβ5.186), (eβ5.191) and (eβ5.197),
| (eβ5.198) | |||||
| (eβ5.199) |
We may write where and are mutually orthogonal non-zero projections in and There are two cases: (1) has real rank zero, then, by Theorem 3.1 of [3], every projection in lifts. If is path connected, by the βMoreoverβ part of Theorem 3.5, we may arrange such that or in Since is unital, Thus, by Lemma 4.1, in both cases, there are mutually orthogonal projections such that There exists such that
| (eβ5.200) |
Then, there are projections such that, with
| (eβ5.201) |
Note that
| (eβ5.202) |
Moreover, by the second part of (eβ5.201) and (eβ5.200), and then, by the first part of (eβ5.201), (eβ5.200), (eβ5.202), we obtain
| (eβ5.203) | |||
| (eβ5.204) | |||
| (eβ5.205) |
Hence
| (eβ5.206) |
It follows from (eβ5.204) and (eβ5.205), and by the assumption on and that
| (eβ5.207) | |||
| (eβ5.208) |
Note Note also that has property (IR). By the choice of and applying Theorem 4.4 of [8], there is a normal element such that
| (eβ5.209) |
Put Then is normal and, by (eβ5.206) and (eβ5.203),
| (eβ5.210) |
Put and Then, since is normal, Moreover
| (eβ5.211) |
β
The Proof of Theorem 1.3.
Proof.
For (i), choose and Note that has real rank zero and stable rank one and has property (IR). So Theorem 5.3 applies.
For (ii), let and If is unital, then As in 4.7, this case follows from Theorem 3.9 (regardless of Fredholm index). Also in the proof of Theorem 1.2 (see 4.7), when is not unital, it is a -unital purely infinite simple -algebraβ which has real rank zero and has property (IR). The fact that is purely infinite is also discussed in 4.7 (the proof of Theorem 1.2). Thus Theorem 5.3 applies.
For (iii), let and As in 4.7 (the proof of Theorem 1.2), by Kasparovβs theorems, and is a -unital hereditary -subalgebraβ of a -unital simple -algebraβ of real rank zero and stable rank one. Hence is also a simple -algebraβ of real rank zero and stable rank one, whence has property (IR). If is unital, then Therefore the conclusion follows from Theorem 4.4 of [8]. Otherwise, by the assumption, has continuous scale (see of [27]). Hence, by Corollary 3.3 of [27], is a simple purely infinite simple -algebra. Thus Theorem 5.3 applies. β
Corollary 5.4.
For there exists satisfying the following: Suppose that is an infinite dimensional separable Hilbert space and with () such that
| (eβ5.212) |
for all Then there are such that
| (eβ5.213) |
Proof.
The next proposition provides a short way to see the second condition in (eβ5.184) is necessary. We also provide a direct proof.
Proposition 5.5.
Let There is satisfying the following: Suppose that and there are such that
| (eβ5.215) |
Then, for any
| (eβ5.216) |
Note that (eβ5.216) also implies that is invertible.
Proof.
Let be given by Proposition 2.15 for (in place of ). Choose Suppose that and are in the statement of the theorem, Put and Then
| (eβ5.217) |
Since is normal, is an essential -near-spectrum of By Proposition 2.15,
| (eβ5.218) |
where By (eβ2.16), we may further have
| (eβ5.219) |
Since is normal, Note, for any
| (eβ5.220) |
Now if then, by (eβ5.219), Since is normal,
| (eβ5.221) |
Hence, by (eβ5.220),
| (eβ5.222) |
This implies that is invertible and
Since we conclude that
β
Remark 5.6.
There is a considerable overlap between the results in this section and that of Theorem 1.1 of [17] even though they have different proofs. However, Theorem 1.1 of [17] provides a quantitative estimate of the distance from to the set of normal elements with finite spectrum (perhaps the quantity is not ready to be computed) while Theorem 5.3 is a conceptional result. In the special case of (i) of Theorem 1.3, (defined in [17]) is equivalent to the condition that for all
On the other hand, computationally speaking, -synthetic-spectrum could be tested. Note that a -synthetic-spectrum consists of finitely many balls and there are only finitely many bounded connected components in its complement. This allows one to avoid the situation of infinitely many βholesβ in the spectrum. Note that the index in each of these βholesβ remains constant. So one only needs to test one point in each of these βholesβ. One also notes that the second condition in (eβ5.184) does not imply
It should also be noted that, when or a normal element ( is Theorem 5.3) may not be approximated by normal elements with finite spectrum even under the assumption that for all
In Theorem 1.1 of [17], seems depending on a given operator The constant in Theorem 5.3 does not depend on nor even -algebras (as long as satisfies the conditions in Theorem 5.3). Finally, let us note that when a unital -algebraβ has (IR), then, by Theorem 4.4 of [8], a pair of almost commuting self-adjoint elements and is close to a pair of commuting ones even though may be large (see also the example below).
Example 5.7.
Let be a unital separable purely infinite simple -algebraβ with and and be a non-unital separable simple -algebraβ with We assume that either is purely infinite simple, or is a separable simple -algebraβ of real rank zero and stable rank one. Consider any unital -algebraβ which is given by the following essential extension:
Note that is purely infinite, has property (IR). Since every projection in lifts to a projection in (see Lemma 4.1). It follows from Corollary 3.16 of [3] that has real rank zero. Hence Theorem 5.3 applies in this case. However, from the six-term exact sequence in -theory, one computes that Denote by the set of normal elements of and the set of those elements in which has finite spectrum, respectively. Then the closure Let be a non-zero projection. Since is either purely infinite or stable rank one, Thus there is a unitary with Let Then is a normal element with Hence for any Since and one sees easily that In other words, Theorem 1.1 in [17] could not be applied to determine when and are almost commuting (see also Theorem in [23]).
6 Epilogue
In this last section, we will address the relation between the condition of vanishing Fredholm index and that of closeness of the (synthetic)-spectrum and essential (synthetic)-spectrum (see Corollary 6.2 and 6.3). We will also give an example (Proposition 6.4) which shows the complexity behind the failure of a higher dimensional version of Theorem 1.3. A final remark on the Mumford original question related to quantum theory and measurements, and their relation with the second condition in Theorem 1.1 is also included.
Let us first state the following corollary of Theorem 4.6.
Corollary 6.1.
Let and There exists satisfying the following: Suppose that with () such that
| (eβ6.223) |
and Then there are such that
| (eβ6.224) |
Proof.
The next corollary shows that, when the second condition in (eβ4.156) is stronger than that of vanishing index in Theorem 1.3.
Corollary 6.2.
For any there exists satisfying the following: Let and be in all cases (i), (ii) and (iii) of Theorem 1.3. Suppose that
| (eβ6.225) |
Then, for any
| (eβ6.226) |
Proof.
Let be given by Proposition 2.16 for
We choose such that the conclusion of Proposition 2.15 holds for Then we choose such that the conclusion of Proposition 2.11 holds for Choose Suppose that (eβ6.225) holds. Put Thus (applying Proposition 2.11) and are all non-empty sets. By Proposition 2.15,
| (eβ6.227) |
By applying Proposition 2.16, we have
| (eβ6.228) |
Applying (eβ2.16), we obtain
| (eβ6.229) |
Since
| (eβ6.230) |
one has (see (eβ2.16))
| (eβ6.231) |
So, if By [5], in case (i) and (ii), since is stable, For case (iii), follows from Lemma 3.3 of [19] (note that, proved in [18], if has real rank zero, then ). Therefore
| (eβ6.232) |
Hence
| (eβ6.233) |
β
In general, we have the following statement. This follows immediately from Theorem 4.6. With the same spirit as that of the proof of Corollary 6.2, one could provide a direct proof without mentioning Theorem 4.6. However, since the point has been made in Corollary 6.2, we refrain ourselves from placing further -theory burden on the reader who might not be familiar with -theory.
Corollary 6.3.
Let For any there exists satisfying the following: Suppose that with () such that
| (eβ6.234) |
where and Then there exists a -brick combination and a unital monomorphism such that
| (eβ6.235) | |||
| (eβ6.236) |
Proof.
Applying Theorem 4.6, one obtains commuting -tuple of self-adjoint operators Let be the -subalgebraβ of generated by Define to be the embedding. Since for any compact metric space we conclude that Consequently where β
We now present an example which shows that an -dimensional version of part (i) of Theorem 1.3 is false (). Note that, in the next proposition, the triple has a single point in the essential -spectrum. Hence the -theory obstruction vanishes in
Proposition 6.4.
There exists and a sequence of triples of self-adjoint operators such that (for all ),
Proof.
Let be the three dimensional unit ball in By Theorem 4.1 of [9], there exists a sequence of contractive linear maps and which satisfy the following:
| (eβ6.237) | |||
| (eβ6.238) |
where the infimum is taken among all homomorphisms
Let and be the quotient map. Then is a homomorphism. Theorem 4.1 of [9] allows each to be -injective (see Definition 0.3 in [9]) with This feature implies that is injective. Let Then there exists a sequence of projections such that Put Note that
| (eβ6.239) |
Hence
| (eβ6.240) |
There are such that, for all sufficiently large
| (eβ6.241) |
Therefore, by replacing with we may assume that maps into and each is unital.
Choose Identify as a finite rank () projection. Put Define by
| (eβ6.242) |
Choose
| (eβ6.243) |
So and Moreover
| (eβ6.244) |
Furthermore,
| (eβ6.245) |
Let and be the quotient map. Define by for
If there is a sequence of commuting triples of self-adjoint operators such that
| (eβ6.246) |
then one obtains a sequence of homomorphisms such that
| (eβ6.247) |
Hence
| (eβ6.248) |
Then where is defined by for
Choose such that if and if Note that and
| (eβ6.249) |
Put and Then
| (eβ6.250) |
Recall that
| (eβ6.251) |
Put and It follows that
| (eβ6.252) |
Let be the -subalgebraβ of generated by Then there is an isomorphism for some compact subset such that and Define by for Since generates we have
| (eβ6.253) |
Let Then,
| (eβ6.254) |
Thus, for all large there exist unitaries such that
| (eβ6.255) |
Define by for Then
| (eβ6.256) |
This contradicts (eβ6.238). Therefore there is no homomorphisms such that (eβ6.248). In other words, there is such that
| (eβ6.257) |
where the infimum is taken among all homomorphisms Combining this with (eβ6.245), the proposition follows. β
Remark 6.5.
Remark 6.6.
This remark was added to the original version of the paper after the author fortunately received a copy of Chapter 14 of a very exciting new book of David Mumford [33] graciously sent by the author. Some terminologies are taken from there.
In quantum theory, one may view macroscopic observables as some self-adjoint operators on a Hilbert space which we assume to be infinite dimensional and separable. Let be the unital -subalgebraβ generated by and (let us assume for convenience here). Suppose that is an orthonormal basis of and suppose that, for any there is such that for all In this case these observables collapse uniformly and any measurements outside (orthogonally) a finite dimensional subspace vanish. In other words, outsiders would see nothing. Let us assume that none of these observables collapse this way (see also Remark 6.5). In other words, we assume that none of βs are compact.
Given a state which is defined by a unit vector the expected value when measuring an observable given by an operator is (see p.178 of [33]). βApproximately Macroscopically Uniqueβ states introduced by Mumford (see Chapter 14 of [33]), or AMU states, are βnear eigenvectorsβ. Set, for a small
where As pointed out in [33], for AMU to be non-empty, it is natural to assume that the commutators have small bound. If the question in the introduction (see (eβ1.1) and (eβ1.2)) had an affirmative answer, then the quantum system could be approximated by a classical one. Consequently, AMU does exist. Observed by Mumford, the condition that the commutators are small is not quite enough (Proposition 6.4 further shows the complexity of the problem) in general.
In an infinite dimensional separable Hilbert space a self-adjoint operator such as may not have any eigenvalues (but have approximate eigenvalues). However, if one allows non-normal states of or that of then does have eigenstates and eigenvalues in the atomic representation of or (see, for example, Proposition 4.3.10 of [34]). In particular, one should consider those states of which vanish on the ideal of compact operators. If a point in the unit cube of is expected to be a βjoint eigenvalueβ, then one may use the test function defined in 2.7 to test it. In order to have (for small ), one only needs one vector state (with a unit vector ) such that
| (eβ6.258) |
One may expect this can be repeatedly measured. Suppose that are unit vectors and is a finite dimensional subspace spanned by Suppose that there is another unit vector such that (eβ6.258) also holds for (in place of ), and this persists. Then we obtain a sequence of () such that (eβ6.258) holds for each (in place of ). Note that unit vector states given by has limit points which give states of vanishing on and Note that there are still unknown (approximate) eigenvalues outside any finite dimensional subspaces of
Let be the quotient map and be a state of the corona algebra (or that of ). Then the quantum state might better reflect this situation. Indeed, one may expect that, at least for one such quantum state one still has
| (eβ6.259) |
Measurements given by states of the form may be called outside measurements. By the Gelfand-Naimark construction, this quantum state is also given by a vector state (the Hilbert space from the atomic representation of or ). In other words,
| (eβ6.260) |
If whenever (eβ6.258) holds, inequality (eβ6.259) holds for some quantum states then the second condition in (eβ1.3) always satisfies (in fact, in this case ).
Let us, for a moment, call a unit vector state for which (eβ6.258) holds a synthetic-AMU state and a synthetic-joint-expected value. The state for which (eβ6.259) holds may be called an essential synthetic-AMU and (for which (eβ6.259) holds) an essential synthetic-joint-expected value. Note that it is easy to see an AMU state (with sufficiently small ) is an ()-synthetic-AMU state. Moreover, by Proposition 2.11, the set of synthetic-AMU states is non-empty, whenever is small (independent of ). The second condition in Theorem 1.1 then may be stated as follows:
(C): all synthetic-joint-expected values are in a sufficiently small neighborhood of the set of essential synthetic-joint-expected values.
This seems to suggest that we could put a condition on the macroscopic system: any local measurement should not be too far off from an outside (or a long-term measurements along an orthonormal basis of the Hilbert space) measurement. Under this condition, when commutators tend to zero, by Theorem 1.1, the quantum system may be approximated by a classical system where observables commute.

### v3

@@
 At the very beginning, he cautioned the audience that there would be no mathematics in his talk.
-Nevertheless, by the very end of this magnificent lecture, when he ended his talk on DNA being a measuring instrument opening a Pandoraβs box, he proposed a problem in -algebras about non-commuting self-adjoint operators to be approximated by commuting ones.
+Nevertheless, by the very end of this magnificent lecture, when he ended his talk on DNA being a measuring instrument opening a Pandora’s box, he proposed a problem in -algebras about non-commuting self-adjoint operators to be approximated by commuting ones.
 A version of this problem proposed by Mumford may be reformulated below: Let and be a positive integer.
@@
 If is any separable Hilbert space and are self-adjoint operators on (with ) such that
-| (eβ1.1) |
+| (e 1.1) |
 there exist self-adjoint operators on such that
-| (eβ1.2) |
+| (e 1.2) |
 In the case that and is any finite dimensional Hilbert space (no bound on the dimension) the exactly the same problem is known as von-Neumann-Kadison-Halmos problem for almost commuting self-adjoint matrices (see [11] and [12]).
 In that case, the problem has an affirmative solution [20] (see also [8] and [13], and, for a fixed dimension, see [32]).
-In infinite dimensional Hilbert space, as observed by Mumford in his new book (p.182), the answer in general is negative (see also Example 4.6 of [8]).
+In an infinite dimensional Hilbert space, as observed by Mumford in his new book (p.182), the answer in general is negative (see also Example 4.6 of [8]) in general.
 The affirmative solution for almost commuting self-adjoint matrices led the further study in the subject of weakly semi-projectivity in -algebras.
 It also has a deep impact on the Elliott program of classification of amenable -algebras (see [7], [24], [26] and [28]).
-The Mumfordβs problem opens a new door to the future applications of this part of -algebraβ theory in previous unexpected areas such as DNA and artificial intelligence ([33]).
+The Mumford’s problem opens a new door to the future applications of this part of -algebra theory in previous unexpected areas such as DNA and artificial intelligence ([34]).
 The first main result of this paper is as follows:
 Theorem 1.1.
-Let and There exists satisfying the following: Suppose that is an infinite dimensional separable Hilbert space and are bounded self-adjoint linear operators on with () such that
-| (eβ1.3) |
+Let and There exists satisfying the following: Suppose that is an infinite dimensional separable Hilbert space and are self-adjoint linear operators on with () such that
+| (e 1.3) |
 where and Then there are bounded self-adjoint linear operators on such that
-| (eβ1.4) |
+| (e 1.4) |
 (See Remark 6.5 for the case that each is compact.)
 Here and are the sets of -synthetic-spectrum, and essential -synthetic-spectrum, respectively (see Definition 2.7).
-In (eβ1.3), is the Hausdorff distance between the two compact subsets of the -dimensional unit cube.
-In the special case that is full, i.e., it contains the whole then the second part of (eβ1.3) is automatically satisfied.
+In (e 1.3), is the Hausdorff distance between the two compact subsets of the -dimensional unit cube.
+In the special case that is full, i.e., it contains the whole then the second part of (e 1.3) is automatically satisfied.
 Remark 6.6 provides some justification of the second condition related to macroscopic observables and measurements, and a discussion of the Mumford problem.
-More generally, let be a Hilbert (right) module over a -algebraβ be the -algebraβ of all bounded module maps on with adjoints and let be the linear span of rank one module maps of the form (for all ), where Denote by the closure of which is a -algebra.
+More generally, let be a Hilbert (right) module over a -algebra be the -algebra of all bounded module maps on with adjoints and let be the linear span of rank one module maps of the form (for all ), where Denote by the closure of which is a -algebra.
 We also have the following:
 Theorem 1.2.
-Let and There exists satisfying the following: Suppose that is a countably generated Hilbert-module over a -unital purely infinite simple -algebraβ and are self-adjoint bounded module maps with () such that
-| (eβ1.5) |
+Let and There exists satisfying the following: Suppose that is a countably generated Hilbert-module over a -unital purely infinite simple -algebra and are self-adjoint bounded module maps with () such that
+| (e 1.5) |
 Suppose also
-(i) if or
-(ii) if and
+(i) or
+(ii) and
 where and
@@
 Then there are such that
-| (eβ1.6) |
+| (e 1.6) |
 For a pair of self-adjoint operators, we also have a more definite answer for some Hilbert modules.
@@
 (i) is an infinite dimensional separable Hilbert space, or
-(ii) is a countably generated Hilbert module over a -unital purely infinite simple -algebraβ
+(ii) is a countably generated Hilbert module over a -unital purely infinite simple -algebra
 such that has real rank zero, or
 (iii) is a countably generated Hilbert module over a -unital simple
--algebraβ of stable rank one such that has real rank zero and is simple.
+-algebra of stable rank one such that has real rank zero and is simple.
 Suppose also that are self-adjoint bounded module maps on with () such that
-| (eβ1.7) |
+| (e 1.7) |
 for all where is the quotient map. Then there are self-adjoint bounded module maps such that
-| (eβ1.8) |
+| (e 1.8) |
 Moreover, the condition that has real rank zero can be removed if we assume
@@
 is connected.
-The second condition in (eβ1.7) means that a Fredholm index vanishes (see Definition 2.4).
+The second condition in (e 1.7) means that a Fredholm index vanishes (see Definition 2.4).
 Note that both case (ii) and case (iii) are generalization of case (i).
@@
 Since it is well-known that spectrum (and the essential spectrum) of a bounded operator is notoriously unstable regarding perturbation, it is necessary to consider some version of -spectrum (which is usually larger than the spectrum).
-For we introduce the notion of -synthetic-spectrum for -tuples of self-adjoint elements in a unital -algebraβ (see Definition 2.7).
-We will also show that, unfortunately, the version of Mumfordβs problem formulated above has a negative solution in general even the associated Fredholm index vanishes.
-In quantum mechanics, one often heard βwhen the commutators go to zero, we recover the classical systemβ.
-If βrecoverβ means the observables are approximated by commuting observables, then example in (iii) of Chapter 14 of [33] (Example 4.6 of [8]), Proposition 5.5 and Proposition 6.4 all show that, mathematically, there are serious (topological) obstacles prevent this type of statements from being true.
+For we introduce the notion of -synthetic-spectrum for -tuples of self-adjoint elements in a unital -algebra (see Definition 2.7).
+We will also show that, unfortunately, the version of Mumford’s problem formulated above has a negative solution in general even the associated Fredholm index vanishes.
+In quantum mechanics, one often heard “when the commutators go to zero, we recover the classical system”.
+If “recover” means the observables are approximated by commuting observables, then example in (iii) of Chapter 14 of [34] (Example 4.6 of [8]), Proposition 5.5 and Proposition 6.4 all show that, mathematically, there are serious (topological) obstacles prevent this type of statements from being true.
 Nevertheless Theorem 1.1 with some interpretation in Remark 6.6 may provide an alternative statement.
@@
 We then present an example that an -tuple version of Theorem 1.3 does not hold in general.
-A final remark was added which contains the justification of the second condition in (eβ1.3) and its connection to the Mumford original question.
+A final remark was added which contains the justification of the second condition in (e 1.3) and its connection to the Mumford original question.
 2 Preliminaries
 Definition 2.1.
-Let be a -algebra.
-Denote by the set of all self-adjoint elements in Denote by the multiplier algebra (the idealizer of in ).
-Denote by and the group of invertible elements and the path connected component of containing the identity of respectively.
-If and we write
-| (eβ2.9) |
+Let be a -algebra. Denote by the set of all self-adjoint elements in Denote by the multiplier algebra (the idealizer of in ). If and we write
+| (e 2.9) |
+For unital -algebra dnote by and the group of invertible elements and the path connected component of containing the identity of respectively.
 Definition 2.2.
-Let be an infinite dimensional separable Hilbert space. Denote by the -algebraβ of all bounded linear operators on and by the -algebraβ of all compact operators on Denote by the quotient map.
+Let be an infinite dimensional separable Hilbert space. Denote by the -algebra of all bounded linear operators on and by the -algebra of all compact operators on Denote by the quotient map.
 Let Recall that the essential spectrum of is the spectrum of
 Definition 2.3.
-Let be a -algebraβ and be a Hilbert (right) module over (see [15]), or Hilbert -module.
-Denote by the -algebraβ of all bounded module maps on with adjoints.
+Let be a -algebra and be a Hilbert (right) module over (see [15]), or Hilbert -module.
+Denote by the -algebra of all bounded module maps on with adjoints.
 A rank one module map is a bounded module map of the form for all (and fixed ), where is the -value inner product on Denote by the linear span of rank one module maps and the (norm) closure of Following Theorem 1 of [15], we identify with
 Definition 2.4.
-Let be a unital -algebraβ and be an invertible element. Define Note that is a unitary. Define
-| (eβ2.10) |
-Then if and only if Let be such that is invertible in Then
-| (eβ2.11) |
+Let be a unital -algebra and be an invertible element. Define Note that is a unitary. Define
+| (e 2.10) |
+Then if and only if when Let be such that is invertible in Then
+| (e 2.11) |
 Definition 2.5.
@@
 Definition 2.7.
-Let In what follows, for each we choose a fixed integer such that Denote We write Then is -dense in Moreover, if
-Let be a unital -algebraβ and be an -tuple of self-adjoint elements in with (). Fix Let be such that if if and is linear in and in Define, for
-| (eβ2.12) | |||||
-| (eβ2.13) |
-Note that we do not assume that mutually commute and the product in (eβ2.13) has a fixed order.
+Let and In what follows, for each we choose a fixed integer such that Denote We write Then is -dense in Moreover, if
+Let be a unital -algebra and be an -tuple of self-adjoint elements in with (). Fix Let be such that if if and is linear in and in Define, for
+| (e 2.12) | |||||
+| (e 2.13) |
+Note that we do not assume that mutually commute and the product in (e 2.13) has a fixed order.
 For write We may write and Set
-| (eβ2.14) |
+| (e 2.14) |
 This union of finitely many closed balls is called -synthetic-spectrum of the -tuple which, by the definition, is compact.
 Let Suppose Then Note that (). Hence It follow that, if then Then, if
-| (eβ2.15) |
+| (e 2.15) |
 Note also that, if then
-| (eβ2.16) |
+| (e 2.16) |
 To see this, let There is with such that There is such that Write and If then Therefore Hence Therefore
+In what follows, we are particularly interested in the case that
 If for some infinite dimensional separable Hilbert space, define
-| (eβ2.17) |
+| (e 2.17) |
 which will be called essential -synthetic-spectrum. It is clear that
-| (eβ2.18) |
+| (e 2.18) |
 Definition 2.8.
@@
 Definition 2.9.
-Let be a unital -algebraβ and for some Let us assume that Fix Suppose that is a compact subset and is a unital c.p.c. map such that
-| (eβ2.19) | |||||
-| (eβ2.20) | |||||
-| (eβ2.21) |
+Let be a unital -algebra and for some Let us assume that Fix Suppose that is a compact subset and is a unital c.p.c. map such that
+| (e 2.19) | |||||
+| (e 2.20) | |||||
+| (e 2.21) |
 for any which has value on an open ball with the center (for some ) and the radius Then we say that is an -near-spectrum of the -tuple We write
@@
 map.
-On the other hand, the -synthetic-spectrum is always defined and in the computation world, it is computable.
+On the other hand, the -synthetic-spectrum is always defined and, in the computation world, it is computable.
 Their relationship is described in Proposition 2.15 below (see also Proposition 2.16).
@@
 Proposition 2.11.
-Fix For any there exists satisfying the following: Suppose that is a unital -algebraβ and with such that
-| (eβ2.22) |
+Fix For any there exists satisfying the following: Suppose that is a unital -algebra and with such that
+| (e 2.22) |
 Then
@@
 Suppose that, for case (1), (or for case (2), ). Then we obtain an and a sequence of unital -algebras and a sequence of -tuples with (), such that
-| (eβ2.23) |
+| (e 2.23) |
 and, (or, for case (2), )
-Consider and quotient map Let Then Let be the -subalgebraβ of generated by Then for some non-empty compact subset of Let be the monomorphism given by with
+Consider and quotient map Let Then Let be the -subalgebra of generated by Then for some non-empty compact subset of Let be the monomorphism given by with
 To show (1), consider Fix Choose such that then Hence, since is a monomorphism,
-| (eβ2.24) |
+| (e 2.24) |
 Recall that Hence there exists such that
-| (eβ2.25) |
+| (e 2.25) |
 It follows that for all In other words, has -synthetic-spectrum containing This proves (1).
 For case (2), we obtain, by the Choi-Effros Theorem ([4]), a c.p.c. map such that Write where each is a c.p.c map, Then
-| (eβ2.26) |
+| (e 2.26) |
 Let be an -dense subset, and with if and if
@@
 Let be a compact metric space and be a finite generating subset.
-Then, for any and any finite subset there exists satisfying the following: Suppose that is a unital -algebraβ and are unital c.p.c. maps such that
-| (eβ2.29) |
+Then, for any and any finite subset there exists satisfying the following: Suppose that is a unital -algebra and are unital c.p.c. maps such that
+| (e 2.29) |
 Then, for all
-| (eβ2.30) |
-Moreover, if and then, for any and and finite subset we may also require that, for all
-| (eβ2.31) |
+| (e 2.30) |
+Moreover, if and then, for any and any finite subset we may also require that, for all
+| (e 2.31) |
 Proof.
 Suppose the lemma is false. Then there exist a finite subset and a sequence of unital -algebras and two sequences of unital c.p.c. maps such that
-| (eβ2.32) | |||
-| (eβ2.33) | |||
-| (eβ2.34) |
+| (e 2.32) | |||
+| (e 2.33) | |||
+| (e 2.34) |
 Moreover, in the case that and there exists and a finite subset such that
-| (eβ2.35) |
+| (e 2.35) |
 for all
-Put Let the quotient map. Then, since is a generating set of by (eβ2.33), is a unital homomorphism, and by (eβ2.32), Hence
-| (eβ2.36) |
+Put Let the quotient map. Then, since is a generating set of by (e 2.33), is a unital homomorphism, and by (e 2.32), Hence
+| (e 2.36) |
 But this implies that, for all sufficiently large
-| (eβ2.37) |
-This leads a contradiction to (eβ2.34). The first part of the lemma then follows.
+| (e 2.37) |
+This leads a contradiction to (e 2.34). The first part of the lemma then follows.
 For the second part, since is a unital homomorphism, for any
-| (eβ2.38) |
+| (e 2.38) |
 It follows that, for all large and
-| (eβ2.39) |
-Another contradiction. So the lemma follows. β
+| (e 2.39) |
+Another contradiction. So the lemma follows. ∎
 Corollary 2.13.
 Fix and There exists satisfying the following: Suppose that is a unital -algebra, with (), is a compact subset, and is a unital c.p.c. map such that
-| (eβ2.40) |
+| (e 2.40) |
 Then, for
-| (eβ2.41) |
+| (e 2.41) |
 Proof.
-Let be as defined in 2.7, where (). There is satisfying the following: for any unital -algebraβ and any pair of self-adjoint elements with if then
-| (eβ2.42) |
+Let be as defined in 2.7, where (). There is satisfying the following: for any unital -algebra and any pair of self-adjoint elements with if then
+| (e 2.42) |
 where is defined in 2.7.
-Let and Choose such that (the second part of) Lemma 2.12 holds for above and (in place of ). Choose such that, for any unital c.p.c. map (any unital -algebraβ ), that
-| (eβ2.43) |
+Let and Choose such that (the second part of) Lemma 2.12 holds for above and (in place of ). Choose such that, for any unital c.p.c. map (any unital -algebra ), that
+| (e 2.43) |
 implies, for
-| (eβ2.44) |
+| (e 2.44) |
 Choose
 Suppose that and satisfy the assumption of this lemma for Since and
-| (eβ2.45) |
-for any -tuple of self-adjoint elements in with (), by the choice of (and ) and by the second part of (eβ2.40) and (eβ2.42), we have, for each
-| (eβ2.46) |
+| (e 2.45) |
+for any -tuple of self-adjoint elements in with (), by the choice of (and ) and by the second part of (e 2.40) and (e 2.42), we have, for each
+| (e 2.46) |
 By the choice of we also have
-| (eβ2.47) |
+| (e 2.47) |
 By the choice and applying (the second part of) Lemma 2.12, we have
-| (eβ2.48) |
-It follows that (keeping in mind of (eβ2.45))
-| (eβ2.49) |
-Combining (eβ2.49) with (eβ2.47), we obtain
-| (eβ2.50) |
-We conclude that, by (eβ2.50) and (eβ2.46),
-| (eβ2.51) | |||
-| (eβ2.52) | |||
-| (eβ2.53) |
-β
+| (e 2.48) |
+It follows that (keeping in mind of (e 2.45))
+| (e 2.49) |
+Combining (e 2.49) with (e 2.47), we obtain
+| (e 2.50) |
+We conclude that, by (e 2.50) and (e 2.46),
+| (e 2.51) | |||
+| (e 2.52) | |||
+| (e 2.53) |
 Definition 2.14.
 Let be a metric space (we only consider the case that ). Recall that the Hausdorff distance of two compact subsets of is defined by
-| (eβ2.54) |
+| (e 2.54) |
 Let be the set of all non-empty compact subsets of Then is a compact metric space with the metric
@@
 Fix For any there exits satisfying the following:
-(1) Suppose that is a unital -algebraβ and with () such that has a non-empty -near-spectrum If is also a non-empty -near-spectrum of and then
-| (eβ2.55) |
+(1) Suppose that is a unital -algebra and with () such that has a non-empty -near-spectrum If is also a non-empty -near-spectrum of and then
+| (e 2.55) |
 (2) Suppose, in addition to (1), has a (closed two-sided) ideal is the quotient map, and and Then
-| (eβ2.56) |
+| (e 2.56) |
 Proof.
@@
 Now suppose that and are -near-spectra of the -tuple Let be c.p.c maps such that
-| (eβ2.57) | |||
-| (eβ2.58) | |||
-| (eβ2.59) |
-for any with for all whenever (if ) or (if ). Let We may assume that Then By (eβ2.59),
-| (eβ2.60) |
-By (eβ2.58) and the choice of applying Lemma 2.12,
-| (eβ2.61) |
+| (e 2.57) | |||
+| (e 2.58) | |||
+| (e 2.59) |
+for any with for all whenever (if ) or (if ). Let We may assume that Then By (e 2.59),
+| (e 2.60) |
+By (e 2.58) and the choice of applying Lemma 2.12,
+| (e 2.61) |
 Since if we have Therefore
-| (eβ2.62) |
+| (e 2.62) |
 This holds for all Exchanging the role of and we obtain
-| (eβ2.63) |
+| (e 2.63) |
 Let and
 By the choice of and applying Corollary 2.13, we have that
-| (eβ2.64) |
+| (e 2.64) |
 Pick There is such that Recall that if As we have Hence for all It follows that
-| (eβ2.65) |
-Thus, by (eβ2.64),
-| (eβ2.66) |
+| (e 2.65) |
+Thus, by (e 2.64),
+| (e 2.66) |
 This implies that
-| (eβ2.67) |
-Next, let By (eβ2.64),
-| (eβ2.68) |
-Hence Thus
-| (eβ2.69) |
-It follows that
-| (eβ2.70) |
-Hence
-| (eβ2.71) |
+| (e 2.67) |
 For (2), let us keep notation above. Then is immediate, and that and that follow from part (1). It remains to show that
 There is a unital c.p.c. map such that
-| (eβ2.72) | |||
-| (eβ2.73) | |||
-| (eβ2.74) |
-for any with if for some Choose We may assume that Hence It follows from (eβ2.74) that
-| (eβ2.75) |
-On the other hand, by the choice of (eβ2.58), (eβ2.73), and applying Lemma 2.12, we have
-| (eβ2.76) |
+| (e 2.72) | |||
+| (e 2.73) | |||
+| (e 2.74) |
+for any with if for some Choose We may assume that Hence It follows from (e 2.74) that
+| (e 2.75) |
+On the other hand, by the choice of (e 2.58), (e 2.73), and applying Lemma 2.12, we have
+| (e 2.76) |
 Thus
-| (eβ2.77) |
+| (e 2.77) |
 Hence It follows that
-| (eβ2.78) |
-Hence β
+| (e 2.78) |
+Hence ∎
 Proposition 2.16.
-Let There exists satisfying the following: Suppose that is a unital -algebraβ and such that (), and
-| (eβ2.79) |
+Let There exists satisfying the following: Suppose that is a unital -algebra and such that (), and
+| (e 2.79) |
 Then
@@
 Otherwise, one obtains a sequence of unital -algebras a sequence of elements with such that
-| (eβ2.80) |
+| (e 2.80) |
 and a sequence of numbers Put and Note that and Since is compact, without loss of generality, we may assume that for some compact subset in as and
 Let and be the quotient map.
-Put Then, by (eβ2.80), is normal.
+Put Then, by (e 2.80), is normal.
 Let Then defined by for is a unital monomorphism.
@@
 Then, since is a homomorphism, for all large is a -near-spectrum for Applying Proposition 2.15, we may assume that, for all
-| (eβ2.83) |
-Note that, by (eβ2.82) and (eβ2.83), Let Note that Hence there are such that ()
-| (eβ2.84) |
+| (e 2.83) |
+Note that, by (e 2.82) and (e 2.83), Let Note that Hence there are such that ()
+| (e 2.84) |
 Hence, for all large
-| (eβ2.85) |
-This is a contradiction. So the lemma follows. β
+| (e 2.85) |
+This is a contradiction. So the lemma follows. ∎
 3 Purely infinite simple -algebras
@@
 An even - brick with corner is a subset of of the form
-| (eβ3.86) |
+| (e 3.86) |
 For the remaining of this paper, all -bricks are even and has the corner An open -brick is the interior of the -brick By boundary of the -brick we mean
@@
 Then
-| (eβ3.87) |
+| (e 3.87) |
 is a pull back. By the Mayer-Vietoris sequence in -theory for -algebras (see [14]), we obtain the following commutative diagram.
-| (eβ3.91) |
+| (e 3.91) |
 Since we have established that and are all finitely generated, we conclude that is also finitely generated. This ends the induction.
-β
 Definition 3.3.
-Let be a compact subset of Define to be the union of all even -bricks (in with corners at ) which has a non-empty intersection with
+Let be a compact subset of Define to be the union of all even -bricks (in with corners at – see Definition 2.6) which has a non-empty intersection with
 We will use the following easy facts:
@@
 We say that has property (F), if is a disjoint union of finitely many path connected compact metric spaces such that is finitely generated (), Choose a base point Write Denote by
-Let be a -algebraβ and a homomorphism. Denote by the restriction.
+Let be a -algebra and a homomorphism. Denote by the restriction.
 Every finite CW complex has property (F). If is a -brick combination in then (by Proposition 3.2) also has property (F).
 Recall that if has finite dimensional range, then there are mutually orthogonal projections and such that
-| (eβ3.92) |
+| (e 3.92) |
 Theorem 3.5.
-Let be a compact metric space with property (F). Suppose that is a unital purely infinite simple -algebraβ and is a unital injective homomorphismβ such that
-| (eβ3.93) |
-Then, for any and any finite subset there exists a set of finitely many mutually orthogonal projections with and such that
-| (eβ3.94) |
+Let be a compact metric space with property (F). Suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism such that
+| (e 3.93) |
+(see Definition 3.4 for ). Then, for any and any finite subset there exists a set of finitely many mutually orthogonal projections with and such that
+| (e 3.94) |
 where for all Moreover, if is connected, we may require that either or in
@@
 Put for some Then we have a splitting short exact sequence
-Recall that is finitely generated (). It follows that, for any unital -algebraβ
-| (eβ3.95) |
-Define by for all Then one computes that, by (eβ3.93),
-| (eβ3.96) |
+Recall that is finitely generated (). It follows that, for any unital -algebra
+| (e 3.95) |
+Define by for all Then one computes that, by (e 3.93),
+| (e 3.96) |
 Fix and finite subset Applying Theorem A of [6], we obtain an integer a unitary and such that, for all
-| (eβ3.97) |
+| (e 3.97) |
 Define by for all Then we may rewrite for all where is a set of mutually orthogonal projections such that Choose mutually orthogonal projections such that where and in (recall that is purely infinite simple).
 Define by for all By replacing by and by we may assume, without loss of generality, that and
-| (eβ3.98) |
+| (e 3.98) |
 (But then is a unitary in ).
-Recall that, by S. Zhangβs theorem ([38]), has real rank zero. Since is injective, by applying Lemma 4.1 of [7], we obtain a set of mutually orthogonal non-zero projections with such that
-| (eβ3.99) |
+Recall that, by S. Zhang’s theorem ([39]), has real rank zero. Since is injective, by applying Lemma 4.1 of [7], we obtain a set of mutually orthogonal non-zero projections with such that
+| (e 3.99) |
 There are such that in (recall again that is purely infinite simple). We may then write
-| (eβ3.100) |
+| (e 3.100) |
 where for all There are such that
-| (eβ3.101) |
+| (e 3.101) |
 Let Then and Moreover,
-| (eβ3.102) |
+| (e 3.102) |
 Hence
-| (eβ3.103) |
-Define for all Then is a unital homomorphismβ from to with finite dimensional range and, by (eβ3.103), (eβ3.100) and (eβ3.98), for all
-Finally, since in as we arranged earlier, by the definition of and the βMoreoverβ part of the statement also holds. β
+| (e 3.103) |
+Define for all Then is a unital homomorphism from to with finite dimensional range and, by (e 3.103), (e 3.100) and (e 3.98), for all
+Finally, since in as we arranged earlier, by the definition of and the “Moreover” part of the statement also holds. ∎
 Let be a compact subset. We use the usual Euclidian metric. Recall that, for any
-| (eβ3.104) |
+| (e 3.104) |
 Theorem 3.6.
 Fix For any and finite subset There exists satisfying the following:
-Let be a compact subset of Suppose that is a unital purely infinite simple -algebraβ and is a unital injective homomorphism.
-If is any compact subset with property (F) such that where then there is a a unital injective homomorphismβ such that
-| (eβ3.105) |
+Let be a compact subset of Suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism.
+If is any compact subset with property (F) such that where then there is a unital injective homomorphism such that
+| (e 3.105) |
 where is the quotient map by restriction ( for all ).
@@
 Let and finite subset be given. Choose such that, for all
-| (eβ3.106) |
-Fix any Suppose that is a compact subset of such that Since is compact, there is with such that
-| (eβ3.107) |
+| (e 3.106) |
+Suppose that is a compact subset of such that Since is compact, there is with such that
+| (e 3.107) |
 Put and Then There is a finite -net such that where Then
@@
 Choose mutually orthogonal non-zero projections such that Define by Define by Then is injective. One computes that
-| (eβ3.108) |
+| (e 3.108) |
 It follows from Theorem 1.7 of [6] that there is a unitary such that
-| (eβ3.109) |
-Applying Theorem 3.5, there is a unital homomorphismβ with finite dimensional range such that
-| (eβ3.110) |
+| (e 3.109) |
+Applying Theorem 3.5, there is a unital homomorphism with finite dimensional range such that
+| (e 3.110) |
 We may write for all where and are mutually orthogonal non-zero projections with By choosing a better approximation if necessary, without loss of generality, we may assume that is -dense in Recall that is -dense in and -dense in For each choose exactly one such that Then, we may assume where is a set of distinct points such that
-| (eβ3.111) |
+| (e 3.111) |
 Note that we view Then set Put where for all
 By the choice of we have, for all
-| (eβ3.112) |
+| (e 3.112) |
 Recall that Hence in There are partial isometries such that and Define Then
-| (eβ3.113) | |||||
-| (eβ3.114) |
+| (e 3.113) | |||||
+| (e 3.114) |
 So is a unitary in Note that
-| (eβ3.115) |
-Then, combining with (eβ3.110) and (eβ3.112), for all
-| (eβ3.116) |
+| (e 3.115) |
+Then, combining with (e 3.110) and (e 3.112), for all
+| (e 3.116) | |||||
 Define by
-| (eβ3.117) |
-Then is a unital injective homomorphism. Moreover, for all by (eβ3.109), the definition of above and by (eβ3.116),
-| (eβ3.118) | |||||
-To see the second formula in (eβ3.105), we may write, from (eβ3.117), since factors through
-| (eβ3.119) |
-β
+| (e 3.117) |
+Then is a unital injective homomorphism. Moreover, for all by (e 3.109), the definition of above and by (e 3.116),
+| (e 3.118) | |||||
+To see the second formula in (e 3.105), we may write, from (e 3.117), since factors through
+| (e 3.119) |
 Corollary 3.7.
 Fix an integer and
-Suppose that is a compact subset of and suppose that is a unital purely infinite simple -algebraβ and is a unital injective homomorphism.
-If is any compact subset with property (F) such that where then there is a unital injective homomorphismβ such that
-| (eβ3.120) |
+Suppose that is a compact subset of and suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism.
+If is any compact subset with property (F) such that then there is a unital injective homomorphism such that
+| (e 3.120) |
 where is the quotient map by restriction ( for all ).
@@
 Suppose that is a unital injective homomorphism.
-Then, for any and any finite subset there exists such that, for any integer with there exists a unital injective homomorphismβ such that
-| (eβ3.122) |
+Then, for any and any finite subset there exists such that, for any integer with there exists a unital injective homomorphism such that
+| (e 3.122) |
 where is the quotient map by restriction, is a finite subset such that (and where is the -brick cover of ).
 Theorem 3.9.
-Let and There is satisfying the following: If is a unital purely infinite simple -algebraβ and with () such that
-| (eβ3.123) |
-then there is integer with and and a -brick combination and a unital injective homomorphismβ such that
-| (eβ3.124) |
+Let and There is satisfying the following: If is a unital purely infinite simple -algebra and with () such that
+| (e 3.123) |
+then there is integer with and and a -brick combination and a unital injective homomorphism such that
+| (e 3.124) |
 Proof.
 Let It follows from Corollary 1.20 of [22] that there exists such that when there are such that
-| (eβ3.125) |
-Let be the commutative -subalgebraβ generated by Then there exists a unital homomorphismβ such that Hence for some compact subset of Denote the isomorphism give above. We may assume that
-Put Choose any with and By Theorem 3.6, we obtain a -brick combination and a unital injective homomorphismβ such that
-| (eβ3.126) |
+| (e 3.125) |
+Let be the commutative -subalgebra generated by Then there exists a unital homomorphism such that Hence for some compact subset of Denote the isomorphism given above. We may assume that
+Put Choose any with and By Theorem 3.6, we obtain a -brick combination and a unital injective homomorphism such that
+| (e 3.126) |
 Hence
-| (eβ3.127) |
-β
+| (e 3.127) |
 4 Multiple self-adjoint operators
@@
 Lemma 4.1.
-Let be a unital -algebraβ and be an ideal of which is a -unital -algebraβ of real rank zero. Suppose that is a projection such that where is the quotient map. Then there is a projection such that
+Let be a unital -algebra and be an ideal of which is a -unital -algebra of real rank zero. Suppose that is a projection such that where is the quotient map. Then there is a projection such that
 Lemma 4.2.
@@
 map where such that
-| (eβ4.128) | |||
-| (eβ4.129) |
+| (e 4.128) | |||
+| (e 4.129) |
 and
 Proof.
-Let be given. and Choose Since is a -near-spectrum of there is a unital c.p.c. map such that
-| (eβ4.130) |
-Since is a -spectrum of there exists a unital monomorphism such that
-| (eβ4.131) |
+Let Choose Since is a -near-spectrum of there is a unital c.p.c. map such that
+| (e 4.130) |
+Since is a -spectrum of there exists, by Corollary 3.7, a compact subset such that and a unital monomorphism such that
+| (e 4.131) |
 Let be a non-zero projection such that Choose a partial isometry such that
-| (eβ4.132) |
-It follows that in Let be a -dense subset of Let be mutually orthogonal non-zero projections in such that in Define by
-| (eβ4.133) |
+| (e 4.132) |
+It follows that in Let be a -dense subset of which is also -dense in Let be mutually orthogonal non-zero projections in such that in Define by
+| (e 4.133) |
 Note that in Then, since is purely infinite and simple, by Theorem 1.7 of [6], there exists a unitary such that
-| (eβ4.134) |
+| (e 4.134) |
 Let and
 Then we may write
-| (eβ4.135) |
+| (e 4.135) |
 Since in by applying Lemma 4.1, we obtain projections such that
-| (eβ4.136) | |||
-| (eβ4.137) |
+| (e 4.136) | |||
+| (e 4.137) |
 Let Then is a projection. By the Choi-Effros lifting theorem ([4]), there is a c.p.c. map such that
@@
 Then is an approximate identity for and is an approximate identity for Choose such that
-| (eβ4.139) |
+| (e 4.139) |
 Hence, since
-| (eβ4.140) |
-Put Then, by (eβ4.138), for
-| (eβ4.142) | |||||
+| (e 4.140) |
+Put Then, by (e 4.138), for
+| (e 4.142) | |||||
 Similarly, for
-| (eβ4.144) | |||||
+| (e 4.144) | |||||
 Put Then
-| (eβ4.145) | |||
-| (eβ4.146) | |||
-| (eβ4.147) |
+| (e 4.145) | |||
+| (e 4.146) | |||
+| (e 4.147) |
 Put for Then it is a c.p.c. map. Moreover
-| (eβ4.149) | |||||
-Hence, using the inequality above and (eβ4.130),
-| (eβ4.150) |
+| (e 4.148) | |||||
+| (e 4.149) |
+Hence, using the inequality above and (e 4.130),
+| (e 4.150) |
 To arrange we may split We write where are two mutually orthogonal projections and both are not in as Then define by for all We then replace with and replace with
@@
 First, by [5], We also have, for any For any projection we have This immediately implies (1), (4) and (7).
-It follows from Theorem 1.1 of [39] that and Therefore, (2), (3), (5), (6), (8) and (9) hold. β
+It follows from Theorem 1.1 of [40] that and Therefore, (2), (3), (5), (6), (8) and (9) hold. ∎
 Definition 4.4.
-Let be a unital -algebra. Let and be maps, and We say satisfies condition (R, r, b, T, s) if has --cancellation with for all -stable rank -divisible rank and for all
+Let be a unital -algebra. Let and be maps, and We say satisfies condition if has --cancellation with for all -stable rank -divisible rank for all and for all
 If is a unital purely infinite simple -algebra, then satisfies condition for any and and If has stable rank one and real rank zero, then satisfies condition for any and
@@
 Theorem 4.5.
-Let where is a compact subset. Let be maps and
-For any there exist a positive number and an integer satisfying the following: Suppose that is a unital -algebraβ satisfying condition such that for any compact metric space and are unital two c.p.c.
+Let where is a compact subset. Let be maps, and
+For any there exist a positive number and an integer satisfying the following: Suppose that is a unital -algebra satisfying condition such that for any compact metric space and are unital c.p.c.
 maps such that
-| (eβ4.151) |
+| (e 4.151) |
 then there are a unitary and a homomorphism with finite dimensional image such that
-| (eβ4.152) |
+| (e 4.152) |
 Proof.
-Let and Since we assume that for any compact metric space for any finite subset as in Theorem 3.1 of [10], we have that for any sufficiently almost multiplicative c.p.c.
+Let and Since we assume that for any compact metric space for any finite subset as in Theorem 3.1 of [10], we have that for any sufficiently multiplicative c.p.c.
 maps Then, by Proposition 4.3 and applying Theorem 3.1 of [10], we obtain and a finite subset such that, if
-| (eβ4.153) |
-there exists an integer a unital homomorphismβ and a unitary such that
-| (eβ4.154) |
+| (e 4.153) |
+there exists an integer a unital homomorphism with finite dimensional range and a unitary such that
+| (e 4.154) |
 (Note as mentioned above, that whenever it makes sense, since for any compact metric space )
 Theorem 4.6.
-Let and Let and be maps and There exists satisfying the following: Suppose that is a unital -algebraβ such that for any compact metric space () satisfying condition and is an essential ideal which has real rank zero such that is purely infinite simple, for any projection and suppose that with () such that
-| (eβ4.156) |
+Let and Let and be maps and There exists satisfying the following: Suppose that is a unital -algebra such that for any compact metric space () satisfying condition and is an essential ideal which has real rank zero such that is purely infinite simple, for any non-zero projection and suppose that with () such that
+| (e 4.156) |
 where and where is the quotient map. Then there are such that
-| (eβ4.157) |
+| (e 4.157) |
 Proof.
@@
 Assume that there are many points in There are at most many compact subsets of such that, for any compact subset there exists such that
-| (eβ4.158) |
+| (e 4.158) |
 We will apply Theorem 4.5. Let (in place of ) be given by Theorem 4.5 for and (in place of ), Define
@@
 Now suppose that with () such that
-| (eβ4.159) |
-By the choice of (and ), by applying Proposition 2.11, and are not empty set (see also (eβ2.15)) and the -tuple has an -near-spectrum and, by Theorem 3.9, has an -spectrum Moreover (by the choice of ) such that (by applying Proposition 2.15),
-| (eβ4.160) |
+| (e 4.159) |
+By the choice of (and ) and, applying Proposition 2.11, and are non-empty set (see also (e 2.15)) and the -tuple has an -near-spectrum and, by Theorem 3.9, has an -spectrum Moreover (by the choice of ), we have (by applying Proposition 2.15),
+| (e 4.160) |
 By the choice of applying Lemma 4.2, we obtain mutually orthogonal nonzero projections a -dense subset a c.p.c. map such that (with )
-| (eβ4.161) | |||
-| (eβ4.162) |
+| (e 4.161) | |||
+| (e 4.162) |
 and Put for all
 Note that, since Choose and define by
-By the choice of and (eβ4.162), as and applying Theorem 4.5, we obtain a unital homomorphismβ with finite dimensional range and a unitary such that
-| (eβ4.163) |
+By the choice of and (e 4.162), as and applying Theorem 4.5, we obtain a unital homomorphism with finite dimensional range and a unitary such that
+| (e 4.163) |
 Put
-| (eβ4.164) |
+| (e 4.164) |
 Since has a finite dimensional range, there are mutually orthogonal projections (for some integer ) such that
-| (eβ4.165) |
+| (e 4.165) |
 where
-By the assumption that and (eβ4.160), is -dense in Note that
-| (eβ4.166) |
+By the assumption that and (e 4.160), is -dense in Note that
+| (e 4.166) |
 Hence is -dense in Let
-| (eβ4.167) | |||||
-| (eβ4.168) |
+| (e 4.167) | |||||
+| (e 4.168) |
 By induction, since is -dense in we obtain such that and where
 Put Define
-| (eβ4.169) |
+| (e 4.169) |
 Then
-| (eβ4.170) |
+| (e 4.170) |
 In we have a partial isometry such that
 Put Then
-| (eβ4.171) |
+| (e 4.171) |
 It follows that
-| (eβ4.172) |
+| (e 4.172) | |||||
+| (e 4.173) |
 Put
-| (eβ4.173) |
-for all Then, by (eβ4.163), (eβ4.164), (eβ4.170), (eβ4.172) and (eβ4.161), for
-| (eβ4.174) |
+| (e 4.174) |
+for all Then, by (e 4.163), (e 4.164), (e 4.170), (e 4.173) and (e 4.161), for
+| (e 4.175) |
 Put
-| (eβ4.175) |
-Then, since by (eβ4.174),
-| (eβ4.176) |
-Since is a homomorphismβ so is Hence
-| (eβ4.177) |
-β
+| (e 4.176) |
+Then, since by (e 4.175),
+| (e 4.177) |
+Since is a homomorphism so is Hence
+| (e 4.178) |
 Proof of Theorem 1.1
-Proof.
 4.7.
@@
 Proof.
-Let be a -unital purely infinite simple -algebraβ and be a countably generated Hilbert -module.
-Let Then, by Kasparovβs absorbing theorem (Theorem 2 of [15]), is an orthogonal summand of and hence is isomorphic to a hereditary -subalgebraβ of which is purely infinite simple.
+Let be a -unital purely infinite simple -algebra and be a countably generated Hilbert -module.
+Let Then, by Kasparov’s absorbing theorem (Theorem 2 of [15]), is an orthogonal summand of and hence is isomorphic to a hereditary -subalgebra of which is purely infinite simple.
 If is unital, by Theorem 1 of [15], Then the conclusion of Theorem 1.2 follows from Theorem 3.9 (without referring to synthetic spectra).
@@
 It follows from Theorem 3.3 of [38] (see also Remark 2.5 (b) of [31]) that is purely infinite simple, and every projection in is equivalent to Thus, by Proposition 4.3, Theorem 4.6 applies.
-β
+∎
 5 A pair of almost commuting self-adjoint operators on Hilbert modules
 Definition 5.1 (Definition 3.1 of [8]).
-For a unital -algebraβ denote by the set of elements with the property that for no ideal of is one-sided and not two-sided invertible in Recall that is said to have property (IR) if all elements in belongs to the norm closure of the group of invertible elements.
-A non-unital -algebraβ has property (IR), if the -algebraβ obtained by adjoining a unit to has property (IR).
+For a unital -algebra denote by the set of elements with the property that for no ideal of is one-sided and not two-sided invertible in Recall that is said to have property (IR) if all elements in belongs to the norm closure of the group of invertible elements.
+A non-unital -algebra has property (IR), if the -algebra obtained by adjoining a unit to has property (IR).
 It follows from Lemma 4 of [23] that, for any projection has (IR).
-Every -algebraβ of stable rank one has property (IR), and every purely infinite simple -algebraβ has property (IR).
+Every -algebra of stable rank one has property (IR), and every purely infinite simple -algebra has property (IR).
 Lemma 5.2.
 Let There exists satisfying the following: Let be a unital purely infinite simple -algebra, be self-adjoint elements such that and
-| (eβ5.178) |
+| (e 5.179) |
 Then there exists a pair of commuting self-adjoint elements such that
-| (eβ5.179) |
+| (e 5.180) |
 whenever
-(Note that the identity in (eβ5.179) also implies that )
+(Note that the identity in (e 5.180) also implies that )
 Proof.
 It follows from Theorem 3.9 that exists such that, whenever
-| (eβ5.180) |
-there exists a compact subset and a unital injective homomorphismβ such that
-| (eβ5.181) |
+| (e 5.181) |
+there exists a compact subset and a unital injective homomorphism such that
+| (e 5.182) |
 Put and Then is a normal element in and Suppose that such that Then
-| (eβ5.182) |
+| (e 5.183) |
 Hence
-| (eβ5.183) |
-It follows that Hence and are in the same path connected component of It follows that β
+| (e 5.184) |
+It follows that Hence and are in the same path connected component of It follows that ∎
 Theorem 5.3.
-Let There exists satisfying the following: Suppose that is a unital -algebraβ with an essential ideal which is a -unital -algebraβ with real rank zero and with property (IR) such that is purely infinite simple.
+Let There exists satisfying the following: Suppose that is a unital -algebra with an essential ideal which is a -unital -algebra with real rank zero and with property (IR) such that is purely infinite simple.
 Suppose that with () such that
-| (eβ5.184) |
+| (e 5.185) |
 for all where is the quotient map. Then, if either has real rank zero, or is path connected, there are such that
-| (eβ5.185) |
+| (e 5.186) |
 Proof.
@@
 Now assume that satisfy the assumption for Put
-Recall that is a purely infinite simple -algebra. Therefore, by the choice of and applying Theorem 5.2 we obtain a compact subset a unital injective homomorphismβ such that
-| (eβ5.186) |
+Recall that is a purely infinite simple -algebra. Therefore, by the choice of and applying Theorem 5.2 we obtain a compact subset a unital injective homomorphism such that
+| (e 5.187) |
 for those for which where and In particular, has an On the other hand, by the choice of we have
-| (eβ5.187) |
+| (e 5.188) |
 We claim: Suppose that is a bounded path connected component and that there is such that Then, for any
-| (eβ5.188) |
-To see the claim, we note that, by (eβ5.187), Then, by (eβ5.186) and by the assumption (eβ5.184),
-| (eβ5.189) |
+| (e 5.189) |
+To see the claim, we note that, by (e 5.188), Then, by (e 5.187) and by the assumption (e 5.185),
+| (e 5.190) |
 It follows that, for any
-| (eβ5.190) |
+| (e 5.191) |
 and the claim follows.
@@
 Put Then
-By Corollary 3.7, there is a unital injective homomorphismβ such that
-| (eβ5.191) |
+By Corollary 3.7, there is a unital injective homomorphism such that
+| (e 5.192) |
 where is the quotient map by restriction. Put and Then
@@
 (ii) if then
-To see this, let us assume it is not the case (i). Then is in a bounded path connected open component ButοΌ since one must have Hence
+To see this, let us assume it is not the case (i). Then is in a bounded path connected open component But, since one must have Hence
 Next we will show that To see this, let where is a bounded path connected component of Put
-| (eβ5.192) |
+| (e 5.193) |
 If (i.e., it is case (i)), then, for
-| (eβ5.193) |
-Hence, by the second part of (eβ5.191),
-| (eβ5.194) |
-If and (i.e., in the case (ii)), by the second part of (eβ5.191) and the claim above, one computes that
-| (eβ5.195) |
+| (e 5.194) |
+Hence, by the second part of (e 5.192),
+| (e 5.195) |
+If and (i.e., in the case (ii)), by the second part of (e 5.192) and the claim above, one computes that
+| (e 5.196) |
 Hence
-| (eβ5.196) |
+| (e 5.197) |
 Since and is finitely generated, is free. Therefore in Then, by Theorem 3.5, there is a normal element with which has finitely many points in the spectrum contained in such that
-| (eβ5.197) |
-Then, by (eβ5.186), (eβ5.191) and (eβ5.197),
-| (eβ5.198) | |||||
-| (eβ5.199) |
+| (e 5.198) |
+Then, by (e 5.187), (e 5.192) and (e 5.198),
+| (e 5.199) | |||||
+| (e 5.200) |
 We may write where and are mutually orthogonal non-zero projections in and There are two cases: (1) has real rank zero, then, by Theorem 3.14 of [3], every projection in lifts.
-If is path connected, by the βMoreoverβ part of Theorem 3.5, we may arrange such that or in Since is unital, Thus, by Lemma 4.1, in both cases, there are mutually orthogonal projections such that There exists such that
-| (eβ5.200) |
+(2) If is path connected, by the “Moreover” part of Theorem 3.5, we may arrange such that or in Since is unital, Thus, by Lemma 4.1, in both cases, there are mutually orthogonal projections such that There exists such that
+| (e 5.201) |
 Then, there are projections such that, with
-| (eβ5.201) |
+| (e 5.202) |
 Note that
-| (eβ5.202) |
-Moreover, by the second part of (eβ5.201) and (eβ5.200), and then, by the first part of (eβ5.201), (eβ5.200), (eβ5.202), we obtain
-| (eβ5.203) | |||
-| (eβ5.204) | |||
-| (eβ5.205) |
+| (e 5.203) |
+Moreover, by the second part of (e 5.202) and (e 5.201), and then, by the first part of (e 5.202), (e 5.201), (e 5.203), we obtain
+| (e 5.204) | |||
+| (e 5.205) | |||
+| (e 5.206) |
 Hence
-| (eβ5.206) |
-It follows from (eβ5.204) and (eβ5.205), and by the assumption on and that
-| (eβ5.207) | |||
-| (eβ5.208) |
+| (e 5.207) |
+It follows from (e 5.205) and (e 5.206), and by the assumption on and that
+| (e 5.208) | |||
+| (e 5.209) |
 Note Note also that has property (IR). By the choice of and applying Theorem 4.4 of [8], there is a normal element such that
-| (eβ5.209) |
-Put Then is normal and, by (eβ5.206) and (eβ5.203),
-| (eβ5.210) |
+| (e 5.210) |
+Put Then is normal and, by (e 5.207) and (e 5.204),
+| (e 5.211) |
 Put and Then, since is normal, Moreover
-| (eβ5.211) |
-β
+| (e 5.212) |
 The Proof of Theorem 1.3.
@@
 For (ii), let and If is unital, then As in 4.7, this case follows from Theorem 3.9 (regardless of Fredholm index).
-Also in the proof of Theorem 1.2 (see 4.7), when is not unital, it is a -unital purely infinite simple -algebraβ which has real rank zero and has property (IR).
+Also in the proof of Theorem 1.2 (see 4.7), when is not unital, it is a -unital purely infinite simple -algebra which has real rank zero and has property (IR).
 The fact that is purely infinite is also discussed in 4.7 (the proof of Theorem 1.2).
 Thus Theorem 5.3 applies.
-For (iii), let and As in 4.7 (the proof of Theorem 1.2), by Kasparovβs theorems, and is a -unital hereditary -subalgebraβ of a -unital simple -algebraβ of real rank zero and stable rank one.
-Hence is also a simple -algebraβ of real rank zero and stable rank one, whence has property (IR).
+For (iii), let and As in 4.7 (the proof of Theorem 1.2), by Kasparov’s theorems, and is a -unital hereditary -subalgebra of a -unital simple -algebra of real rank zero and stable rank one.
+Hence is also a simple -algebra of real rank zero and stable rank one, whence has property (IR).
 If is unital, then Therefore the conclusion follows from Theorem 4.4 of [8].
@@
 Thus Theorem 5.3 applies.
-β
+∎
 Corollary 5.4.
 For there exists satisfying the following: Suppose that is an infinite dimensional separable Hilbert space and with () such that
-| (eβ5.212) |
+| (e 5.213) |
 for all Then there are such that
-| (eβ5.213) |
+| (e 5.214) |
 Proof.
-The next proposition provides a short way to see the second condition in (eβ5.184) is necessary. We also provide a direct proof.
+The next proposition provides a short way to see the second condition in (e 5.185) is necessary. We also provide a direct proof.
 Proposition 5.5.
 Let There is satisfying the following: Suppose that and there are such that
-| (eβ5.215) |
+| (e 5.216) |
 Then, for any
-| (eβ5.216) |
-Note that (eβ5.216) also implies that is invertible.
+| (e 5.217) |
+Note that (e 5.217) also implies that is invertible.
 Proof.
 Let be given by Proposition 2.15 for (in place of ). Choose Suppose that and are in the statement of the theorem, Put and Then
-| (eβ5.217) |
+| (e 5.218) |
 Since is normal, is an essential -near-spectrum of By Proposition 2.15,
-| (eβ5.218) |
-where By (eβ2.16), we may further have
-| (eβ5.219) |
+| (e 5.219) |
+where By (e 2.16), we may further have
+| (e 5.220) |
 Since is normal, Note, for any
-| (eβ5.220) |
-Now if then, by (eβ5.219), Since is normal,
-| (eβ5.221) |
-Hence, by (eβ5.220),
-| (eβ5.222) |
+| (e 5.221) |
+Now if then, by (e 5.220), Since is normal,
+| (e 5.222) |
+Hence, by (e 5.221),
+| (e 5.223) |
 This implies that is invertible and
 Since we conclude that
-β
 Remark 5.6.
@@
 Note that a -synthetic-spectrum consists of finitely many balls and there are only finitely many bounded connected components in its complement.
-This allows one to avoid the situation of infinitely many βholesβ in the spectrum.
-Note that the index in each of these βholesβ remains constant.
-So one only needs to test one point in each of these βholesβ.
-One also notes that the second condition in (eβ5.184) does not imply
+This allows one to avoid the situation of infinitely many “holes” in the spectrum.
+Note that the index in each of these “holes” remains constant.
+So one only needs to test one point in each of these “holes”.
+One also notes that the second condition in (e 5.185) does not imply
 It should also be noted that, when or a normal element ( is Theorem 5.3) may not be approximated by normal elements with finite spectrum even under the assumption that for all
 In Theorem 1.1 of [17], seems depending on a given operator The constant in Theorem 5.3 does not depend on nor even -algebras (as long as satisfies the conditions in Theorem 5.3).
-Finally, let us note that when a unital -algebraβ has (IR), then, by Theorem 4.4 of [8], a pair of almost commuting self-adjoint elements and is close to a pair of commuting ones even though may be large (see also the example below).
+Finally, let us note that when a unital -algebra has (IR), then, by Theorem 4.4 of [8], a pair of almost commuting self-adjoint elements and is close to a pair of commuting ones even though may be large (see also the example below).
 Example 5.7.
-Let be a unital separable purely infinite simple -algebraβ with and and be a non-unital separable simple -algebraβ with We assume that either is purely infinite simple, or is a separable simple -algebraβ of real rank zero and stable rank one.
-Consider any unital -algebraβ which is given by the following essential extension:
+Let be a unital separable purely infinite simple -algebra with and and be a non-unital separable simple -algebra with We assume that either is purely infinite simple, or is a separable simple -algebra of real rank zero and stable rank one.
+Consider any unital -algebra which is given by the following essential extension:
 Note that is purely infinite, has property (IR).
@@
 Let and There exists satisfying the following: Suppose that with () such that
-| (eβ6.223) |
+| (e 6.224) |
 and Then there are such that
-| (eβ6.224) |
+| (e 6.225) |
 Proof.
-The next corollary shows that, when the second condition in (eβ4.156) is stronger than that of vanishing index in Theorem 1.3.
+The next corollary shows that, when the second condition in (e 4.156) is stronger than that of vanishing index in Theorem 1.3.
 Corollary 6.2.
 For any there exists satisfying the following: Let and be in all cases (i), (ii) and (iii) of Theorem 1.3. Suppose that
-| (eβ6.225) |
+| (e 6.226) |
 Then, for any
-| (eβ6.226) |
+| (e 6.227) |
 Proof.
 Let be given by Proposition 2.16 for
-We choose such that the conclusion of Proposition 2.15 holds for Then we choose such that the conclusion of Proposition 2.11 holds for Choose Suppose that (eβ6.225) holds.
+We choose such that the conclusion of Proposition 2.15 holds for Then we choose such that the conclusion of Proposition 2.11 holds for Choose Suppose that (e 6.226) holds.
 Put Thus (applying Proposition 2.11) and are all non-empty sets.
 By Proposition 2.15,
-| (eβ6.227) |
+| (e 6.228) |
 By applying Proposition 2.16, we have
-| (eβ6.228) |
-Applying (eβ2.16), we obtain
-| (eβ6.229) |
+| (e 6.229) |
+Applying (e 2.16), we obtain
+| (e 6.230) |
 Since
-| (eβ6.230) |
-one has (see (eβ2.16))
-| (eβ6.231) |
+| (e 6.231) |
+one has (see (e 2.16))
+| (e 6.232) |
 So, if By [5], in case (i) and (ii), since is stable, For case (iii), follows from Lemma 3.3 of [19] (note that, proved in [18], if has real rank zero, then ). Therefore
-| (eβ6.232) |
+| (e 6.233) |
 Hence
-| (eβ6.233) |
-β
+| (e 6.234) |
 In general, we have the following statement.
@@
 Let For any there exists satisfying the following: Suppose that with () such that
-| (eβ6.234) |
+| (e 6.235) |
 where and Then there exists a -brick combination and a unital monomorphism such that
-| (eβ6.235) | |||
-| (eβ6.236) |
+| (e 6.236) | |||
+| (e 6.237) |
 Proof.
-Applying Theorem 4.6, one obtains commuting -tuple of self-adjoint operators Let be the -subalgebraβ of generated by Define to be the embedding.
-Since for any compact metric space we conclude that Consequently where β
+Applying Theorem 4.6, one obtains commuting -tuple of self-adjoint operators Let be the -subalgebra of generated by Define to be the embedding.
+Since for any compact metric space we conclude that Consequently where ∎
 We now present an example which shows that an -dimensional version of part (i) of Theorem 1.3 is false ().
@@
 Let be the three dimensional unit ball in By Theorem 4.1 of [9], there exists a sequence of contractive linear maps and which satisfy the following:
-| (eβ6.237) | |||
-| (eβ6.238) |
+| (e 6.238) | |||
+| (e 6.239) |
 where the infimum is taken among all homomorphisms
@@
 Let Then there exists a sequence of projections such that Put Note that
-| (eβ6.239) |
+| (e 6.240) |
 Hence
-| (eβ6.240) |
+| (e 6.241) |
 There are such that, for all sufficiently large
-| (eβ6.241) |
+| (e 6.242) |
 Therefore, by replacing with we may assume that maps into and each is unital.
 Choose Identify as a finite rank () projection. Put Define by
-| (eβ6.242) |
+| (e 6.243) |
 Choose
-| (eβ6.243) |
+| (e 6.244) |
 So and Moreover
-| (eβ6.244) |
+| (e 6.245) |
 Furthermore,
-| (eβ6.245) |
+| (e 6.246) |
 Let and be the quotient map. Define by for
 If there is a sequence of commuting triples of self-adjoint operators such that
-| (eβ6.246) |
-then one obtains a sequence of homomorphisms such that
-| (eβ6.247) |
+| (e 6.247) |
+then one obtains a sequence of unital homomorphisms such that
+| (e 6.248) |
 Hence
-| (eβ6.248) |
+| (e 6.249) |
 Then where is defined by for
 Choose such that if and if Note that and
-| (eβ6.249) |
+| (e 6.250) |
 Put and Then
-| (eβ6.250) |
+| (e 6.251) |
 Recall that
-| (eβ6.251) |
+| (e 6.252) |
 Put and It follows that
-| (eβ6.252) |
-Let be the -subalgebraβ of generated by Then there is an isomorphism for some compact subset such that and Define by for Since generates we have
-| (eβ6.253) |
+| (e 6.253) |
+Let be the (commutative) -subalgebra of generated by Then there is an isomorphism for some compact subset such that and Define by for Since generates we have
+| (e 6.254) |
 Let Then,
-| (eβ6.254) |
+| (e 6.255) |
 Thus, for all large there exist unitaries such that
-| (eβ6.255) |
+| (e 6.256) |
 Define by for Then
-| (eβ6.256) |
-This contradicts (eβ6.238). Therefore there is no homomorphisms such that (eβ6.248). In other words, there is such that
-| (eβ6.257) |
-where the infimum is taken among all homomorphisms Combining this with (eβ6.245), the proposition follows. β
+| (e 6.257) |
+This contradicts (e 6.239). Therefore there is no homomorphisms such that (e 6.249). In other words, there is such that
+| (e 6.258) |
+where the infimum is taken among all homomorphisms Combining this with (e 6.246), the proposition follows. ∎
 Remark 6.5.
@@
 In quantum theory, one may view macroscopic observables as some self-adjoint operators on a Hilbert space which we assume to be infinite dimensional and separable.
-Let be the unital -subalgebraβ generated by and (let us assume for convenience here).
+Let be the unital -subalgebra generated by and (let us assume for convenience here).
 Suppose that is an orthonormal basis of and suppose that, for any there is such that for all In this case these observables collapse uniformly and any measurements outside (orthogonally) a finite dimensional subspace vanish.
@@
 Let us assume that none of these observables collapse this way (see also Remark 6.5).
-In other words, we assume that none of βs are compact.
+In other words, we assume that none of ’s are compact.
 Given a state which is defined by a unit vector the expected value when measuring an observable given by an operator is (see p.178 of [34]).
-βApproximately Macroscopically Uniqueβ states introduced by Mumford (see Chapter 14 of [33]), or AMU states, are βnear eigenvectorsβ.
+“Approximately Macroscopically Unique” states introduced by Mumford (see Chapter 14 of [34]), or AMU states, are “near eigenvectors”.
 Set, for a small
 where As pointed out in [34], for AMU to be non-empty, it is natural to assume that the commutators have small bound.
-If the question in the introduction (see (eβ1.1) and (eβ1.2)) had an affirmative answer, then the quantum system could be approximated by a classical one.
+If the question in the introduction (see (e 1.1) and (e 1.2)) had an affirmative answer, then the quantum system could be approximated by a classical one.
 Consequently, AMU does exist.
@@
 In particular, one should consider those states of which vanish on the ideal of compact operators.
-If a point in the unit cube of is expected to be a βjoint eigenvalueβ, then one may use the test function defined in 2.7 to test it.
+If a point in the unit cube of is expected to be a “joint eigenvalue”, then one may use the test function defined in 2.7 to test it.
 In order to have (for small ), one only needs one vector state (with a unit vector ) such that
-| (eβ6.258) |
+| (e 6.259) |
 One may expect this can be repeatedly measured.
-Suppose that are unit vectors and is a finite dimensional subspace spanned by Suppose that there is another unit vector such that (eβ6.258) also holds for (in place of ), and this persists.
-Then we obtain a sequence of () such that (eβ6.258) holds for each (in place of ).
+Suppose that are unit vectors and is a finite dimensional subspace spanned by Suppose that there is another unit vector such that (e 6.259) also holds for (in place of ), and this persists.
+Then we obtain a sequence of () such that (e 6.259) holds for each (in place of ).
 Note that unit vector states given by has limit points which give states of vanishing on and Note that there are still unknown (approximate) eigenvalues outside any finite dimensional subspaces of
@@
 Indeed, one may expect that, at least for one such quantum state one still has
-| (eβ6.259) |
+| (e 6.260) |
 Measurements given by states of the form may be called outside measurements.
@@
 In other words,
-| (eβ6.260) |
-If whenever (eβ6.258) holds, inequality (eβ6.259) holds for some quantum states then the second condition in (eβ1.3) always satisfies (in fact, in this case ).
-Let us, for a moment, call a unit vector state for which (eβ6.258) holds a synthetic-AMU state and a synthetic-joint-expected value.
-The state for which (eβ6.259) holds may be called an essential synthetic-AMU and (for which (eβ6.259) holds) an essential synthetic-joint-expected value.
+| (e 6.261) |
+If whenever (e 6.259) holds, inequality (e 6.260) holds for some quantum states then the second condition in (e 1.3) always satisfies (in fact, in this case ).
+Let us, for a moment, call a unit vector state for which (e 6.259) holds a synthetic-AMU state and a synthetic-joint-expected value.
+The state for which (e 6.260) holds may be called an essential synthetic-AMU and (for which (e 6.260) holds) an essential synthetic-joint-expected value.
 Note that it is easy to see an AMU state (with sufficiently small ) is an ()-synthetic-AMU state.

### v4

@@
 We study the problem when an -tuple of self-adjoint operators in an infinite dimensional separable Hilbert space with small commutators is close to an -tuple of commuting self-adjoint operators on We give an affirmative answer to the problem when the synthetic-spectrum and the essential synthetic-spectrum are close.
-In the case that we show that a pair of almost commuting self-adjoint operators in an infinite dimensional separable Hilbert space is close to a commuting pair of self-adjoint operators if and only if a corresponding Fredholm index vanishes outside of an essential synthetic-spectrum.
 Examples are also exhibited that, in general, the answer to the problem when is negative even the associated Fredholm index vanishes.
@@
 1 Introduction
-In the First International Congress of Basic Science held in 2023, David Mumford gave the opening plenary lecture on Consciousness, robots and DNA ([33]).
-At the very beginning, he cautioned the audience that there would be no mathematics in his talk.
-Nevertheless, by the very end of this magnificent lecture, when he ended his talk on DNA being a measuring instrument opening a Pandora’s box, he proposed a problem in -algebras about non-commuting self-adjoint operators to be approximated by commuting ones.
-A version of this problem proposed by Mumford may be reformulated below: Let and be a positive integer.
-Is there a constant such that the following statement holds?
-If is any separable Hilbert space and are self-adjoint operators on (with ) such that
+At the First International Congress of Basic Science held in 2023, David Mumford delivered the opening plenary lecture on Consciousness, Robots and DNA ([33]).
+He began by cautioning the audience that there would be no mathematics in his talk.
+Nevertheless, by the very end of this magnificent lecture—which culminated with the idea of DNA as a measuring instrument opening a Pandora’s box— he proposed an unexpected problem in -algebras: determining when almost-commuting self-adjoint operators can be approximated by commuting ones.
+A version of this problem proposed by Mumford may be reformulated as follows: Let and be a positive integer.
+When is there a constant such that the following statement holds?
+For any separable Hilbert space and self-adjoint operators (with ) satisfying
 | (e 1.1) |
@@
 In the case that and is any finite dimensional Hilbert space (no bound on the dimension) the exactly the same problem is known as von-Neumann-Kadison-Halmos problem for almost commuting self-adjoint matrices (see [11] and [12]).
-In that case, the problem has an affirmative solution [20] (see also [8] and [13], and, for a fixed dimension, see [32]).
-In an infinite dimensional Hilbert space, as observed by Mumford in his new book (p.182), the answer in general is negative (see also Example 4.6 of [8]) in general.
-The affirmative solution for almost commuting self-adjoint matrices led the further study in the subject of weakly semi-projectivity in -algebras.
+In this setting, the problem has an affirmative solution [20] (see also [8] and [13], and [32] for fixed dimensions).
+However, in an infinite dimensional Hilbert space as Mumford noted in his recent book (p.182) – and as was previously known – the answer is negative in general (see also Example 4.6 of [8]).
+The affirmative solution for almost commuting self-adjoint matrices has spurred further research in the study of weak semi-projectivity in -algebras.
 It also has a deep impact on the Elliott program of classification of amenable -algebras (see [7], [24], [26] and [28]).
-The Mumford’s problem opens a new door to the future applications of this part of -algebra theory in previous unexpected areas such as DNA and artificial intelligence ([34]).
+Moreover, Mumford’s problem opens new avenues for applying this branch of -algebra theory to previously unexpected fields, such as DNA research and artificial intelligence ([34]).
 The first main result of this paper is as follows:
@@
 | (e 1.6) |
-For a pair of self-adjoint operators, we also have a more definite answer for some Hilbert modules.
+While we are mostly interested in the case that we also looks the special case when For a pair of self-adjoint operators, we have a more definite answer for some Hilbert modules.
 Theorem 1.3.
@@
 The second condition in (e 1.7) means that a Fredholm index vanishes (see Definition 2.4).
-Note that both case (ii) and case (iii) are generalization of case (i).
-In case (ii), is stable (whenever is not a finitely generated projective Hilbert module), and is simple (like the Calkin algebra).
-In case (ii), we keep the feature that is simple and has real rank zero.
-Case (i) in Theorem 1.3 is a more definite result than that of Theorem 1.1 in the case We actually prove a slightly more general result (see Theorem 5.3 below).
+Note that both cases (ii) and (iii) generalize case (i).
+In case (ii), is stable (whenever is not a finitely generated projective Hilbert module), and is simple (analogous to the Calkin algebra).
+Moreover, in case (ii), retains its simiplicity while has real rank zero.
+Case (i) in Theorem 1.3 is a more definite result than that of Theorem 1.1 for In fact, we prove a slightly more general result (see Theorem 5.3 below).
 One should also compare (i) in Theorem 1.3 to Theorem 1.1 of [17] (see Remark 5.6).
-In the case that however, the normal elements may not be approximated by normal elements in with finite spectrum.
-We also show that condition of the vanishing index associated with the pair of operators is necessary.
-Since it is well-known that spectrum (and the essential spectrum) of a bounded operator is notoriously unstable regarding perturbation, it is necessary to consider some version of -spectrum (which is usually larger than the spectrum).
-For we introduce the notion of -synthetic-spectrum for -tuples of self-adjoint elements in a unital -algebra (see Definition 2.7).
-We will also show that, unfortunately, the version of Mumford’s problem formulated above has a negative solution in general even the associated Fredholm index vanishes.
-In quantum mechanics, one often heard “when the commutators go to zero, we recover the classical system”.
-If “recover” means the observables are approximated by commuting observables, then example in (iii) of Chapter 14 of [34] (Example 4.6 of [8]), Proposition 5.5 and Proposition 6.4 all show that, mathematically, there are serious (topological) obstacles prevent this type of statements from being true.
-Nevertheless Theorem 1.1 with some interpretation in Remark 6.6 may provide an alternative statement.
+Theorem 1.1 of [17] provides a more definitive estimate of the distance from an element to the set of normal elements with finite spectrum in -algebras with real rank zero.
+However, when normal elements may not be approximated by normal elements in with finite spectrum.
+Thus, cases (ii) and (iii) in Theorem 1.3 cover scenarios where almost normal elements are close to normal elements that cannot be approximated normal elements with finite spectra — complementing the results in [17].
+We further demonstrate that the vanishing index condition for the operator pair is necessary.
+Given well-known instablity of the spectrum (and essential spectrum) of bounded operators under perturbations, it becomes essential to consider some version of -spectrum (typically larger than the spectrum).
+For we introduce the concept of -synthetic-spectrum for -tuples of self-adjoint elements in a unital -algebra (see Definition 2.7).
+We will also show that, unfortunately, the version of Mumford’s problem formulated above has a negative solution in general— even the associated Fredholm index vanishes when
+In quantum mechanics, one often encouters the statement, “when the commutators go to zero, we recover the classical system”.
+If “recover” means the observables are approximated by commuting observables, then example in (iii) of Chapter 14 of [34] (Example 4.6 of [8]), Proposition 5.5 and Proposition 6.4 all reveal significnat topological obstructions to such claims.
+Nevertheless Theorem 1.1, interpreted as in Remark 6.6, may offer an alternative perpective.
 The paper is organized as follows.
@@
 Definition 2.3.
-Let be a -algebra and be a Hilbert (right) module over (see [15]), or Hilbert -module.
+Let be a -algebra and a Hilbert (right) module over (or Hilbert -module; see [15]).
 Denote by the -algebra of all bounded module maps on with adjoints.
-A rank one module map is a bounded module map of the form for all (and fixed ), where is the -value inner product on Denote by the linear span of rank one module maps and the (norm) closure of Following Theorem 1 of [15], we identify with
+A rank-one module map is a bounded module map of the form for all (and fixed ), where is the -valued inner product on Denote by the linear span of rank-one module maps and the (norm) closure of Following Theorem 1 of [15], we identify with
 Definition 2.4.
@@
 In by we mean the euclidian distance between and
+i.e.,
 Denote by If and define
@@
 Definition 2.6.
-Fix an integer Define
-has only finitely many points.
+Fix Fix an integer Define
+has only finitely many points. Most of the time, we choose
 Definition 2.7.
@@
 Note that we do not assume that mutually commute and the product in (e 2.13) has a fixed order.
-For write We may write and Set
+For write We may sometime write and Set
 | (e 2.14) |
@@
 | (e 2.16) |
-To see this, let There is with such that There is such that Write and If then Therefore Hence Therefore
+To see this, let There is with such that There is such that Write and If then
+Therefore Hence Therefore
 In what follows, we are particularly interested in the case that
@@
 | (e 2.21) |
-for any which has value on an open ball with the center (for some ) and the radius Then we say that is an -near-spectrum of the -tuple We write
-If, moreover, is a unital homomorphism, then we say is an - spectrum of the -tuple
-Let be an -near-spectrum and be a -near -spectrum for respectively. Suppose that then, by the definition, is also an -near-spectrum. In particular, if then
+for any which has value on an open ball with the center (for some ) and the radius Then we say that is an -near-spectrum of the -tuple We may write
+for convenince. The reader is recommended to read Remark 2.10 for clarification.
+If is, additionally, a unital homomorphism, then we say is an - spectrum of the -tuple
 Let be an infinite dimensional separable Hilbert space and let If is an -spectrum of we say that is an essential -spectrum for In this case we write
 Remark 2.10.
-We often use the -near spectrum as it associated with a c.p.c.
-map, whenever the -tuple sufficiently almost commutative as shown in the next proposition (2.11).
-But it is not unique and depends on the choice of the c.p.c.
+Here are some clarifications.
+1.
+It should be noted that is not uniquely determined and depends on the choice of the c.p.c.
 map.
-On the other hand, the -synthetic-spectrum is always defined and, in the computation world, it is computable.
+Writing simply means that there exists a non-empty compact subset and a unital c.p.c.
+map satisfying conditions (i), (ii) and (iii) of Definition 2.9.
+This convention is adopted for notational convenience.
+Moreover, if we write we mean that no such and exist which satisfy (i), (ii) and (iii).
+2. However, we will later show that -near spectra are unique up to an - neighborhood when the -tuple selfadjoint operators are almost commuting within as elaborated in Proposition 2.15.
+3.
+On the other hand, the -synthetic-spectrum is always uniquely defined and—importantly– computable in the context of computations.
 Their relationship is described in Proposition 2.15 below (see also Proposition 2.16).
-In practice, it is quite possible that it is more difficult to compute the spectrum of an operator where than to compute the norm of
+4. Let be an -near-spectrum and be a -near -spectrum for respectively. Suppose that then, by the definition, is also an -near-spectrum. In particular, if then
+5. In practice, computing the spectrum of an operator (where ) may be more challenging than computing the norm of
 Proposition 2.11.
@@
 Then
-(1) and
+(1) and,
+(2) i.e., there exists a non-empty compact subset and a c.p.c. map satisfies condition (i), (ii) and (iii) in Definition 2.9.
 Proof.
-Suppose that, for case (1), (or for case (2), ). Then we obtain an and a sequence of unital -algebras and a sequence of -tuples with (), such that
+Suppose that the lemma is false. Then we obtain an and a sequence of unital -algebras and a sequence of -tuples with (), such that
 | (e 2.23) |
-and, (or, for case (2), )
+and, for case (1), (or, for case (2), )
 Consider and quotient map Let Then Let be the -subalgebra of generated by Then for some non-empty compact subset of Let be the monomorphism given by with
-To show (1), consider Fix Choose such that then Hence, since is a monomorphism,
+To show (1), consider (see Definition 2.7). Fix Choose such that then Hence, since is a monomorphism,
 | (e 2.24) |
-Recall that Hence there exists such that
+Recall that Hence there is an infinite subset such that, for all
 | (e 2.25) |
-It follows that for all In other words, has -synthetic-spectrum containing This proves (1).
-For case (2), we obtain, by the Choi-Effros Theorem ([4]), a c.p.c. map such that Write where each is a c.p.c map, Then
+It follows that for all In other words, has -synthetic-spectrum containing for all This contradicts “ ”. Thus (1) follows.
+For (2), we obtain, by the Choi-Effros Theorem ([4]), a c.p.c. map such that Write where each is a c.p.c map, Then
 | (e 2.26) |
@@
 where is defined in 2.7.
-Let and Choose such that (the second part of) Lemma 2.12 holds for above and (in place of ). Choose such that, for any unital c.p.c. map (any unital -algebra ), that
+Let and For choose such that (the second part of) Lemma 2.12 holds for above and (in place of ). Choose such that, for any unital c.p.c. map (any unital -algebra ), that
 | (e 2.43) |
@@
 | (e 2.47) |
-By the choice and applying (the second part of) Lemma 2.12, we have
+Define by for all By the choice and applying (the second part of) Lemma 2.12 to we obtain
 | (e 2.48) |
-It follows that (keeping in mind of (e 2.45))
+In other words,
 | (e 2.49) |
-Combining (e 2.49) with (e 2.47), we obtain
+Note that It follows that (keeping in mind of (e 2.45))
 | (e 2.50) |
-We conclude that, by (e 2.50) and (e 2.46),
-| (e 2.51) | |||
+Combining (e 2.50) with (e 2.47), we obtain
+| (e 2.51) |
+We conclude that, by (e 2.51) and (e 2.46),
 | (e 2.52) | |||
-| (e 2.53) |
+| (e 2.53) | |||
+| (e 2.54) |
 Definition 2.14.
-Let be a metric space (we only consider the case that ). Recall that the Hausdorff distance of two compact subsets of is defined by
-| (e 2.54) |
-Let be the set of all non-empty compact subsets of Then is a compact metric space with the metric
+Let be a metric space (we only consider the case that is compact such as the case that ). Recall that the Hausdorff distance of two compact subsets of is defined by
+| (e 2.55) |
+Let be the set of all non-empty compact subsets of Then is a compact metric space with the metric when is compact.
 Proposition 2.15.
 Fix For any there exits satisfying the following:
-(1) Suppose that is a unital -algebra and with () such that has a non-empty -near-spectrum If is also a non-empty -near-spectrum of and then
-| (e 2.55) |
+(1) Suppose that is a unital -algebra and with () such that has a non-empty -near-spectrum If is also a non-empty -near-spectrum of then and
+| (e 2.56) |
 (2) Suppose, in addition to (1), has a (closed two-sided) ideal is the quotient map, and and Then
-| (e 2.56) |
+| (e 2.57) |
 Proof.
@@
 Let be an -dense subset of Choose with if and if Put
-Note that is a generating set of Let be given by Lemma 2.12 (for and ) and be given by Corollary 2.13. Choose
-Now suppose that and are -near-spectra of the -tuple Let be c.p.c maps such that
-| (e 2.57) | |||
+Note that is a generating set of Let be given by Lemma 2.12 (for and as well as ) and be given by Corollary 2.13. Choose
+Now suppose that and are -near-spectra of the -tuple Let be unital c.p.c maps such that
 | (e 2.58) | |||
-| (e 2.59) |
-for any with for all whenever (if ) or (if ). Let We may assume that Then By (e 2.59),
-| (e 2.60) |
-By (e 2.58) and the choice of applying Lemma 2.12,
+| (e 2.59) | |||
+| (e 2.60) | |||
 | (e 2.61) |
-Since if we have Therefore
+for any with for all whenever (for ), or (for ).
+Let We may assume that Then By (e 2.61),
 | (e 2.62) |
+Define () by and for all By (e 2.59) and applying Lemma 2.12, we have, for
+| (e 2.63) |
+Then, by (e 2.62), we obtain, for
+| (e 2.64) |
+Since if we must have Therefore
+| (e 2.65) |
 This holds for all Exchanging the role of and we obtain
-| (e 2.63) |
+| (e 2.66) |
 Let and
 By the choice of and applying Corollary 2.13, we have that
-| (e 2.64) |
+| (e 2.67) |
 Pick There is such that Recall that if As we have Hence for all It follows that
-| (e 2.65) |
-Thus, by (e 2.64),
-| (e 2.66) |
+| (e 2.68) |
+Thus, by (e 2.67),
+| (e 2.69) |
 This implies that
-| (e 2.67) |
-For (2), let us keep notation above. Then is immediate, and that and that follow from part (1). It remains to show that
+| (e 2.70) |
+Since we conclude that The same argument shows that
+Next, let By (e 2.67),
+| (e 2.71) |
+Hence Thus
+| (e 2.72) |
+It follows that
+| (e 2.73) |
+Hence
+| (e 2.74) |
+This completes the proof of part (1).
+For (2), let us keep notation above. Then is immediate.
+Suppose that Then that follow from part (1). It remains to show that
 There is a unital c.p.c. map such that
-| (e 2.72) | |||
-| (e 2.73) | |||
-| (e 2.74) |
-for any with if for some Choose We may assume that Hence It follows from (e 2.74) that
-| (e 2.75) |
-On the other hand, by the choice of (e 2.58), (e 2.73), and applying Lemma 2.12, we have
-| (e 2.76) |
-Thus
+| (e 2.75) | |||
+| (e 2.76) | |||
 | (e 2.77) |
-Hence It follows that
-| (e 2.78) |
-Hence ∎
+for any with if for some
 Proposition 2.16.
 Let There exists satisfying the following: Suppose that is a unital -algebra and such that (), and
-| (e 2.79) |
+| (e 2.83) |
 Then
@@
 Otherwise, one obtains a sequence of unital -algebras a sequence of elements with such that
-| (e 2.80) |
-and a sequence of numbers Put and Note that and Since is compact, without loss of generality, we may assume that for some compact subset in as and
+| (e 2.84) |
+and a sequence of numbers Put and Note that and Since is compact, without loss of generality, by passing to a subsequence, we may assume that for some compact subset in as and
 Let and be the quotient map.
-Put Then, by (e 2.80), is normal.
+Put Then, by (e 2.84), is normal.
 Let Then defined by for is a unital monomorphism.
@@
 maps such that Let be as in Proposition 2.15.
-Then, since is a homomorphism, for all large is a -near-spectrum for Applying Proposition 2.15, we may assume that, for all
-| (e 2.83) |
-Note that, by (e 2.82) and (e 2.83), Let Note that Hence there are such that ()
-| (e 2.84) |
+Then, since is a homomorphism, for all large is a (non-empty) -near-spectrum for Applying Proposition 2.15, we may assume that, for all
+| (e 2.87) |
+Note that, by (e 2.86) and (e 2.87), Let Note that Hence there are such that ()
+| (e 2.88) |
 Hence, for all large
-| (e 2.85) |
+| (e 2.89) |
 This is a contradiction. So the lemma follows. ∎
@@
 Definition 3.1.
-An even - brick with corner is a subset of of the form
-| (e 3.86) |
-For the remaining of this paper, all -bricks are even and has the corner An open -brick is the interior of the -brick By boundary of the -brick we mean
-A -brick combination is a finite union of -bricks. The -brick combination has the following easy property: If are two distinct -bricks of then
-We think Proposition 3.2 are well known.
+An even - brick with corner (with – see Definition 2.6) is a subset of of the form
+| (e 3.90) |
+For the remaining of this paper, all -bricks are assumed to be even and have their corner An open -brick is the interior of the -brick The boundary of a -brick is defined as
+A -brick combination is a finite union of -bricks. Since two distinct - bricks have distinct corners, a -brick combination has the following easy property: If are two distinct -bricks of then
+We believe that Proposition 3.2 to be well known.
 Proposition 3.2.
@@
 We prove this by induction on the number of convex sets.
-If is convex, is contractive. Hence and
+If is convex, then is contractive. Hence and
 Suppose that the proposition is proved for many closed convex sets,
@@
 Then
-| (e 3.87) |
+| (e 3.91) |
 is a pull back. By the Mayer-Vietoris sequence in -theory for -algebras (see [14]), we obtain the following commutative diagram.
-| (e 3.91) |
-Since we have established that and are all finitely generated, we conclude that is also finitely generated. This ends the induction.
+| (e 3.95) |
+Since we have established that and are all finitely generated abelian groups, we conclude that is also finitely generated. This ends the induction.
 Definition 3.3.
@@
 Recall that if has finite dimensional range, then there are mutually orthogonal projections and such that
-| (e 3.92) |
+| (e 3.96) |
 Theorem 3.5.
 Let be a compact metric space with property (F). Suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism such that
-| (e 3.93) |
+| (e 3.97) |
 (see Definition 3.4 for ). Then, for any and any finite subset there exists a set of finitely many mutually orthogonal projections with and such that
-| (e 3.94) |
+| (e 3.98) |
 where for all Moreover, if is connected, we may require that either or in
 Proof.
-Since has property (F), we may write where each is path connected.
-Let be such that and if Put ().
+Since has property (F), we may write where each is path connected and is finitely generated (), Let be such that and if Put ().
 By considering for each we reduce the general case to the case that has only one path connected component.
-Put for some Then we have a splitting short exact sequence
+Note that, if originally, is not connected, then we do not need to consider the “Moreover” part of the theorem.
+If is connected, then we need to point out that
+So now we assume that is connected. Put for some Then we have a splitting short exact sequence
 Recall that is finitely generated (). It follows that, for any unital -algebra
-| (e 3.95) |
-Define by for all Then one computes that, by (e 3.93),
-| (e 3.96) |
+| (e 3.99) |
+Define by for all Then one computes that, by (e 3.97),
+| (e 3.100) |
 Fix and finite subset Applying Theorem A of [6], we obtain an integer a unitary and such that, for all
-| (e 3.97) |
+| (e 3.101) |
 Define by for all Then we may rewrite for all where is a set of mutually orthogonal projections such that Choose mutually orthogonal projections such that where and in (recall that is purely infinite simple).
 Define by for all By replacing by and by we may assume, without loss of generality, that and
-| (e 3.98) |
+| (e 3.102) |
 (But then is a unitary in ).
 Recall that, by S. Zhang’s theorem ([39]), has real rank zero. Since is injective, by applying Lemma 4.1 of [7], we obtain a set of mutually orthogonal non-zero projections with such that
-| (e 3.99) |
+| (e 3.103) |
 There are such that in (recall again that is purely infinite simple). We may then write
-| (e 3.100) |
+| (e 3.104) |
 where for all There are such that
-| (e 3.101) |
+| (e 3.105) |
 Let Then and Moreover,
-| (e 3.102) |
+| (e 3.106) |
 Hence
-| (e 3.103) |
-Define for all Then is a unital homomorphism from to with finite dimensional range and, by (e 3.103), (e 3.100) and (e 3.98), for all
-Finally, since in as we arranged earlier, by the definition of and the “Moreover” part of the statement also holds. ∎
+| (e 3.107) |
+Define, for all
+| (e 3.108) |
+Then is a unital homomorphism from to with finite dimensional range and, by (e 3.107), (e 3.104) and (e 3.102), for all
+Finally, since in as we arranged earlier, by (e 3.108), the “Moreover” part of the statement also holds. ∎
 Let be a compact subset. We use the usual Euclidian metric. Recall that, for any
-| (e 3.104) |
+| (e 3.109) |
 Theorem 3.6.
@@
 Let be a compact subset of Suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism.
-If is any compact subset with property (F) such that where then there is a unital injective homomorphism such that
-| (e 3.105) |
+If is any compact subset with property (F) such that then there is a unital injective homomorphism such that
+| (e 3.110) |
 where is the quotient map by restriction ( for all ).
@@
 Let and finite subset be given. Choose such that, for all
-| (e 3.106) |
-Suppose that is a compact subset of such that Since is compact, there is with such that
-| (e 3.107) |
-Put and Then There is a finite -net such that where Then
-Choose a non-zero projection with Since is a purely infinite simple -algebra, there is a unital embedding Applying Theorem 2.8 of [16], one obtains a unital embedding Note in There is a partial isometry such that and Define by for all Then is a unital injective homomorphism.
-Choose mutually orthogonal non-zero projections such that Define by Define by Then is injective. One computes that
-| (e 3.108) |
+| (e 3.111) |
+Suppose that is a compact subset of such that which has property (F). Since is compact, there is with such that
+| (e 3.112) |
+Put and Then There is a finite -net such that where Then Note that we assume that if
+Choose a non-zero projection with in and Since is a purely infinite simple -algebra, there is a unital embedding Applying Theorem 2.8 of [16], one obtains a unital embedding Note in There is a partial isometry such that and Define by for all Then is a unital injective homomorphism.
+Choose mutually orthogonal non-zero projections such that Define by Then in Define by for all Then is injective. One computes that
+| (e 3.113) |
 It follows from Theorem 1.7 of [6] that there is a unitary such that
-| (e 3.109) |
+| (e 3.114) |
 Applying Theorem 3.5, there is a unital homomorphism with finite dimensional range such that
-| (e 3.110) |
-We may write for all where and are mutually orthogonal non-zero projections with By choosing a better approximation if necessary, without loss of generality, we may assume that is -dense in Recall that is -dense in and -dense in For each choose exactly one such that Then, we may assume where is a set of distinct points such that
-| (e 3.111) |
+| (e 3.115) |
+We may write for all where and are mutually orthogonal non-zero projections in with By choosing a better approximation if necessary, without loss of generality, we may assume that is -dense in Recall that is -dense in and -dense in For each choose exactly one such that Then, we may assume where is a set of distinct points such that
+| (e 3.116) |
 Note that we view Then set Put where for all
 By the choice of we have, for all
-| (e 3.112) |
+| (e 3.117) |
 Recall that Hence in There are partial isometries such that and Define Then
-| (e 3.113) | |||||
-| (e 3.114) |
+| (e 3.118) | |||||
+| (e 3.119) |
 So is a unitary in Note that
-| (e 3.115) |
-Then, combining with (e 3.110) and (e 3.112), for all
-| (e 3.116) | |||||
+| (e 3.120) |
+Then, combining with (e 3.115) and (e 3.117), for all
+| (e 3.121) | |||||
 Define by
-| (e 3.117) |
-Then is a unital injective homomorphism. Moreover, for all by (e 3.109), the definition of above and by (e 3.116),
-| (e 3.118) | |||||
-To see the second formula in (e 3.105), we may write, from (e 3.117), since factors through
-| (e 3.119) |
+| (e 3.122) |
+Then is a unital injective homomorphism. Moreover, for all by (e 3.114), the definition of and above and by (e 3.121),
+| (e 3.123) | |||||
+To see the second formula in (e 3.110), we may write, from (e 3.122), since factors through
+| (e 3.124) |
 Corollary 3.7.
 Fix an integer and
-Suppose that is a compact subset of and suppose that is a unital purely infinite simple -algebra and is a unital injective homomorphism.
+Suppose that is a compact subset of is a unital purely infinite simple -algebra and is a unital injective homomorphism.
 If is any compact subset with property (F) such that then there is a unital injective homomorphism such that
-| (e 3.120) |
+| (e 3.125) |
 where is the quotient map by restriction ( for all ).
@@
 Corollary 3.8.
-Let be a compact metric space which is a a compact subset of (for integer ) and be a unital purely infinite simple -algebra.
-Suppose that is a unital injective homomorphism.
-Then, for any and any finite subset there exists such that, for any integer with there exists a unital injective homomorphism such that
-| (e 3.122) |
-where is the quotient map by restriction, is a finite subset such that (and where is the -brick cover of ).
+Let Then, for any and any finite subset there exists satisfying the following: Suppose that is a purely infinite simple -algebra and is a unital injective homomorphism, where is a compact subset.
+For any integer with there exists a unital injective homomorphism such that
+| (e 3.127) |
+where is the quotient map by restriction, is a finite subset such that (and where is a -brick cover of ).
 Theorem 3.9.
 Let and There is satisfying the following: If is a unital purely infinite simple -algebra and with () such that
-| (e 3.123) |
+| (e 3.128) |
 then there is integer with and and a -brick combination and a unital injective homomorphism such that
-| (e 3.124) |
+| (e 3.129) |
 Proof.
 Let It follows from Corollary 1.20 of [22] that there exists such that when there are such that
-| (e 3.125) |
+| (e 3.130) |
 Let be the commutative -subalgebra generated by Then there exists a unital homomorphism such that Hence for some compact subset of Denote the isomorphism given above. We may assume that
-Put Choose any with and By Theorem 3.6, we obtain a -brick combination and a unital injective homomorphism such that
-| (e 3.126) |
+Put Choose any with and By Corollary 3.8, we obtain a -brick combination and a unital injective homomorphism such that
+| (e 3.131) |
 Hence
-| (e 3.127) |
+| (e 3.132) |
 4 Multiple self-adjoint operators
@@
 Lemma 4.2.
-Fix For any there exists satisfying the following: Suppose that is a unital -algebra, is a separable ideal of real rank zero, is the quotient map such that is purely infinite simple, and with () such that the -tuple has a -near-spectrum and has -spectrum Then there are mutually orthogonal non-zero projections a -dense subset in and a c.p.c.
-map where such that
-| (e 4.128) | |||
-| (e 4.129) |
+Fix For any there exists satisfying the following: Suppose that is a unital -algebra with an ideal such that:
+1. is separable and has real rank zero;
+2. the quotient is purely infinite ans simple.
+3. with ();
+4. the -tuple has a -near-spectrum and has a -spectrum Then, there are mutually orthogonal non-zero projections a -dense subset in and a c.p.c. map where such that
+| (e 4.133) | |||
+| (e 4.134) |
 and
 Proof.
-Let Choose Since is a -near-spectrum of there is a unital c.p.c. map such that
-| (e 4.130) |
-Since is a -spectrum of there exists, by Corollary 3.7, a compact subset such that and a unital monomorphism such that
-| (e 4.131) |
+Let Choose Since is a -near-spectrum of by Definition 2.9, there is a unital c.p.c. map such that
+| (e 4.135) |
+Since is a -spectrum of there exists, by Definition 2.9 and by Corollary 3.7, a compact subset such that and a unital monomorphism such that
+| (e 4.136) |
 Let be a non-zero projection such that Choose a partial isometry such that
-| (e 4.132) |
+| (e 4.137) |
 It follows that in Let be a -dense subset of which is also -dense in Let be mutually orthogonal non-zero projections in such that in Define by
-| (e 4.133) |
+| (e 4.138) |
 Note that in Then, since is purely infinite and simple, by Theorem 1.7 of [6], there exists a unitary such that
-| (e 4.134) |
+| (e 4.139) |
 Let and
 Then we may write
-| (e 4.135) |
+| (e 4.140) |
 Since in by applying Lemma 4.1, we obtain projections such that
-| (e 4.136) | |||
-| (e 4.137) |
+| (e 4.141) | |||
+| (e 4.142) |
 Let Then is a projection. By the Choi-Effros lifting theorem ([4]), there is a c.p.c. map such that
@@
 Then is an approximate identity for and is an approximate identity for Choose such that
-| (e 4.139) |
+| (e 4.144) |
 Hence, since
-| (e 4.140) |
-Put Then, by (e 4.138), for
-| (e 4.142) | |||||
+| (e 4.145) |
+Put Then, by (e 4.143), for
+| (e 4.147) | |||||
 Similarly, for
-| (e 4.144) | |||||
+| (e 4.149) | |||||
 Put Then
-| (e 4.145) | |||
-| (e 4.146) | |||
-| (e 4.147) |
+| (e 4.150) | |||
+| (e 4.151) | |||
+| (e 4.152) |
 Put for Then it is a c.p.c. map. Moreover
-| (e 4.148) | |||||
-| (e 4.149) |
-Hence, using the inequality above and (e 4.130),
-| (e 4.150) |
+| (e 4.154) | |||||
+Hence, using the inequality above and (e 4.135),
+| (e 4.155) |
 To arrange we may split We write where are two mutually orthogonal projections and both are not in as Then define by for all We then replace with and replace with
@@
 Definition 4.4.
-Let be a unital -algebra. Let and be maps, and We say satisfies condition if has --cancellation with for all -stable rank -divisible rank for all and for all
+Let be a unital -algebra. Let and be maps, and We say satisfies condition if has --cancellation, -stable rank -divisible rank for all and for all
 If is a unital purely infinite simple -algebra, then satisfies condition for any and and If has stable rank one and real rank zero, then satisfies condition for any and
-We will apply the following theorem which is a variation of that of Theorem 3.1 of [10].
+We will apply the following theorem which is a variation of Theorem 3.1 of [10].
 Theorem 4.5.
-Let where is a compact subset. Let be maps, and
-For any there exist a positive number and an integer satisfying the following: Suppose that is a unital -algebra satisfying condition such that for any compact metric space and are unital c.p.c.
-maps such that
-| (e 4.151) |
-then there are a unitary and a homomorphism with finite dimensional image such that
-| (e 4.152) |
-Proof.
-Let and Since we assume that for any compact metric space for any finite subset as in Theorem 3.1 of [10], we have that for any sufficiently multiplicative c.p.c.
-maps Then, by Proposition 4.3 and applying Theorem 3.1 of [10], we obtain and a finite subset such that, if
-| (e 4.153) |
-there exists an integer a unital homomorphism with finite dimensional range and a unitary such that
-| (e 4.154) |
-(Note as mentioned above, that whenever it makes sense, since for any compact metric space )
-Theorem 4.6.
-Let and Let and be maps and There exists satisfying the following: Suppose that is a unital -algebra such that for any compact metric space () satisfying condition and is an essential ideal which has real rank zero such that is purely infinite simple, for any non-zero projection and suppose that with () such that
+Fix Let be maps, and
+For any there exist a positive number and an integer satisfying the following: Suppose that is a unital -algebra such that:
+1. satisfies condition ;
+2. for any compact metric space
+Let be a compact subset of and let be unital c.p.c. maps such that
 | (e 4.156) |
-where and where is the quotient map. Then there are such that
+Then, there are a unitary (for some ) and a homomorphism with finite dimensional image, such that
 | (e 4.157) |
 Proof.
-Fix Let Fix a finite -dense subset such as for some large (as in 2.6).
-Assume that there are many points in There are at most many compact subsets of such that, for any compact subset there exists such that
+Let us first fix a compact subset Let and Since we assume that for any compact metric space for any finite subset as in Theorem 3.1 of [10], we have that for any sufficiently multiplicative c.p.c.
+maps Then, by Proposition 4.3 and applying Theorem 3.1 of [10], we obtain and a finite subset such that, if
 | (e 4.158) |
-We will apply Theorem 4.5. Let (in place of ) be given by Theorem 4.5 for and (in place of ), Define
+there exists an integer a unital homomorphism with finite dimensional range and a unitary such that
+| (e 4.159) |
+(Note as mentioned above, that whenever it makes sense, since for any compact metric space )
+However, since generates as -algebra, one obtains such that when (e 4.156) holds, then
+| (e 4.160) |
+So Theorem 3.1 of [10] applies and the theorem would follow if we fix first.
+To find a common and indepedent of we note that there are finitely many compact subsets of such that, for any non-empty compact subset there is such that
+| (e 4.161) |
+(see 2.14). Let Then, for any non-empty compact subset of there is such that
+| (e 4.162) |
+When is given, by the first part of the proof, we obtain and such that the conclusion of the first part holds for Choose and
+Now suppose and are unital c.p.c. maps satisfying the assumption (e 4.156). Suppose that
+| (e 4.163) |
+Define by (this makes sense as ), Therefore
+| (e 4.164) |
+and Hence, with the choice of we obtain homomorphism with finite dimensional range and a unitary such that, for
+| (e 4.165) |
+We may write where () and is a set of mutually orthogonal projections in Since for each we may choose such that
+| (e 4.166) |
+Define by
+| (e 4.167) |
+Note that
+| (e 4.168) |
+Hence, by (e 4.165), for
+| (e 4.169) |
+Theorem 4.6.
+Let and Let and be maps and There exists satisfying the following:
+Suppose that is a unital -algebra such that:
+1. satisfies condition
+2. for any compact metric space ();
+3. is an essential ideal which has real rank zero such that is purely infinite simple;
+4. for any and any non-zero projection we have
+Suppose further that with () satisfy :
+| (e 4.170) |
+(recall Definition 2.14), where and where is the quotient map.
+Then there are such that
+| (e 4.171) |
+Proof.
 Put Let (in place of ) be given by Lemma 4.2 for (in place of ). Choose Put
@@
 Now suppose that with () such that
-| (e 4.159) |
-By the choice of (and ) and, applying Proposition 2.11, and are non-empty set (see also (e 2.15)) and the -tuple has an -near-spectrum and, by Theorem 3.9, has an -spectrum Moreover (by the choice of ), we have (by applying Proposition 2.15),
-| (e 4.160) |
-By the choice of applying Lemma 4.2, we obtain mutually orthogonal nonzero projections a -dense subset a c.p.c. map such that (with )
-| (e 4.161) | |||
-| (e 4.162) |
-and Put for all
+| (e 4.172) |
+By the choice of (and ) and, applying Proposition 2.11, we obtain non-empty sets and (see also (e 2.15)), and the -tuple has an -near-spectrum and, by Theorem 3.9, has an -spectrum Moreover (by the choice of ), since we have (by applying Proposition 2.15),
+| (e 4.173) |
+By the choice of (and ), applying Lemma 4.2, we obtain mutually orthogonal nonzero projections a -dense subset a c.p.c. map such that (with )
+| (e 4.174) | |||
+| (e 4.175) |
+and Put
 Note that, since Choose and define by
-By the choice of and (e 4.162), as and applying Theorem 4.5, we obtain a unital homomorphism with finite dimensional range and a unitary such that
-| (e 4.163) |
+By the choice of and (e 4.175), as and applying Theorem 4.5, we obtain a unital homomorphism (for some integer ) with finite dimensional range and a unitary such that
+| (e 4.176) |
+Since has a finite dimensional range, there are mutually orthogonal projections (for some integer ) such that
+| (e 4.177) |
+where Put Note that, without loss of generality, by replacing by and adding a projection to each if neseccarily, we may assume that
+Define by
+| (e 4.178) |
+Let be a chosen matrix unit for In what follows, we identify with and above with where In particular, we view
+Recall that is -dense in By the assumption that and (e 4.173), we conclude that is -dense in Note that
+| (e 4.179) |
+Hence is -dense in Let
+| (e 4.180) | |||||
+| (e 4.181) |
+By induction, since is -dense in we obtain mutually disjoint subsets such that and where
+Put Define
+| (e 4.182) |
+Then
+| (e 4.183) |
+Recall that, for any nonzero projection Thus, in viewing we have a partial isometry such that
+Then put Then
+| (e 4.184) |
+It follows that
+| (e 4.185) | |||||
+| (e 4.186) |
 Put
-| (e 4.164) |
-Since has a finite dimensional range, there are mutually orthogonal projections (for some integer ) such that
-| (e 4.165) |
-where
-By the assumption that and (e 4.160), is -dense in Note that
-| (e 4.166) |
-Hence is -dense in Let
-| (e 4.167) | |||||
-| (e 4.168) |
-By induction, since is -dense in we obtain such that and where
-Put Define
-| (e 4.169) |
-Then
-| (e 4.170) |
-In we have a partial isometry such that
-Put Then
-| (e 4.171) |
-It follows that
-| (e 4.172) | |||||
-| (e 4.173) |
+| (e 4.187) |
+for all Then, by (e 4.176), (e 4.178), (e 4.183), (e 4.186) and (e 4.174), for
+| (e 4.188) |
 Put
-| (e 4.174) |
-for all Then, by (e 4.163), (e 4.164), (e 4.170), (e 4.173) and (e 4.161), for
-| (e 4.175) |
-Put
-| (e 4.176) |
-Then, since by (e 4.175),
-| (e 4.177) |
+| (e 4.189) |
+Then, since by (e 4.188),
+| (e 4.190) |
 Since is a homomorphism so is Hence
-| (e 4.178) |
+| (e 4.191) |
 Proof of Theorem 1.1
@@
 A non-unital -algebra has property (IR), if the -algebra obtained by adjoining a unit to has property (IR).
-It follows from Lemma 4 of [23] that, for any projection has (IR).
+It follows from Lemma 4 of [23] that, for nonzero any projection has (IR).
 Every -algebra of stable rank one has property (IR), and every purely infinite simple -algebra has property (IR).
@@
 Let There exists satisfying the following: Let be a unital purely infinite simple -algebra, be self-adjoint elements such that and
-| (e 5.179) |
+| (e 5.192) |
 Then there exists a pair of commuting self-adjoint elements such that
-| (e 5.180) |
-whenever
-(Note that the identity in (e 5.180) also implies that )
+| (e 5.193) |
+(recall Definition 2.4), whenever
+(Note that the identity in (e 5.193) also implies that )
 Proof.
 It follows from Theorem 3.9 that exists such that, whenever
-| (e 5.181) |
+| (e 5.194) |
 there exists a compact subset and a unital injective homomorphism such that
-| (e 5.182) |
-Put and Then is a normal element in and Suppose that such that Then
-| (e 5.183) |
-Hence
-| (e 5.184) |
+| (e 5.195) |
+Put and Then is a normal element in and Moreover
+| (e 5.196) |
+Suppose that such that Then
+| (e 5.197) |
+Hence (also by (e 5.196))
+| (e 5.198) |
 It follows that Hence and are in the same path connected component of It follows that ∎
 Theorem 5.3.
-Let There exists satisfying the following: Suppose that is a unital -algebra with an essential ideal which is a -unital -algebra with real rank zero and with property (IR) such that is purely infinite simple.
-Suppose that with () such that
-| (e 5.185) |
-for all where is the quotient map. Then, if either has real rank zero, or is path connected, there are such that
-| (e 5.186) |
+Let There exists satisfying the following: Suppose that is a unital -algebra with an essential ideal such that:
+1. is -unital, has real rank zero and satisfies property (IR);
+2. the quotient is purely infinite simple.
+Let with () such that:
+| (e 5.199) |
+for all where is the quotient map.
+Then, if either:
+has real rank zero, or
+is path connected,
+there are such that
+| (e 5.200) |
 Proof.
 Fix Let be given by Theorem 4.4 of [8] for in place of
-Choose Let be given by Proposition 2.15 for (in place of ) and Choose Let be given by Proposition 2.16 for (in place of ). Let be given by Theorem 5.2 for (in place of ).
+Choose Let be given by Proposition 2.15 for (in place of ) and Choose
+Let be given by Theorem 5.2 for (in place of ). Put
+Let be given by Proposition 2.16 for (in place of ).
 Choose
-Now assume that satisfy the assumption for Put
+Now assume that satisfy the assumption for By Proposition 2.16, Put (see (e 2.15))
 Recall that is a purely infinite simple -algebra. Therefore, by the choice of and applying Theorem 5.2 we obtain a compact subset a unital injective homomorphism such that
-| (e 5.187) |
-for those for which where and In particular, has an On the other hand, by the choice of we have
-| (e 5.188) |
+| (e 5.201) |
+for those for which where and In particular, the pair has an -spectrum We may view it as an On the other hand, by the choice of and applying Proposition 2.15, we have
+| (e 5.202) |
 We claim: Suppose that is a bounded path connected component and that there is such that Then, for any
-| (e 5.189) |
-To see the claim, we note that, by (e 5.188), Then, by (e 5.187) and by the assumption (e 5.185),
-| (e 5.190) |
+| (e 5.203) |
+To see the claim, we note that, by (e 5.202), Then, by (e 5.201) and by the assumption (e 5.199),
+| (e 5.204) |
 It follows that, for any
-| (e 5.191) |
+| (e 5.205) |
 and the claim follows.
@@
 By Corollary 3.7, there is a unital injective homomorphism such that
-| (e 5.192) |
+| (e 5.206) |
 where is the quotient map by restriction. Put and Then
@@
 Next we will show that To see this, let where is a bounded path connected component of Put
-| (e 5.193) |
+| (e 5.207) |
 If (i.e., it is case (i)), then, for
-| (e 5.194) |
-Hence, by the second part of (e 5.192),
-| (e 5.195) |
-If and (i.e., in the case (ii)), by the second part of (e 5.192) and the claim above, one computes that
-| (e 5.196) |
+| (e 5.208) |
+Hence, by the second part of (e 5.206),
+| (e 5.209) |
+If and (i.e., in the case (ii)), by the second part of (e 5.206) and the claim above, one computes that
+| (e 5.210) |
 Hence
-| (e 5.197) |
+| (e 5.211) |
 Since and is finitely generated, is free. Therefore in Then, by Theorem 3.5, there is a normal element with which has finitely many points in the spectrum contained in such that
-| (e 5.198) |
-Then, by (e 5.187), (e 5.192) and (e 5.198),
-| (e 5.199) | |||||
-| (e 5.200) |
+| (e 5.212) |
+Then, by (e 5.201), (e 5.206) and (e 5.212),
+| (e 5.213) | |||||
+| (e 5.214) |
 We may write where and are mutually orthogonal non-zero projections in and There are two cases: (1) has real rank zero, then, by Theorem 3.14 of [3], every projection in lifts.
 (2) If is path connected, by the “Moreover” part of Theorem 3.5, we may arrange such that or in Since is unital, Thus, by Lemma 4.1, in both cases, there are mutually orthogonal projections such that There exists such that
-| (e 5.201) |
+| (e 5.215) |
 Then, there are projections such that, with
-| (e 5.202) |
+| (e 5.216) |
 Note that
-| (e 5.203) |
-Moreover, by the second part of (e 5.202) and (e 5.201), and then, by the first part of (e 5.202), (e 5.201), (e 5.203), we obtain
-| (e 5.204) | |||
-| (e 5.205) | |||
-| (e 5.206) |
+| (e 5.217) |
+Moreover, by the second part of (e 5.216) and (e 5.215), and then, by the first part of (e 5.216), (e 5.215), (e 5.217), we obtain
+| (e 5.218) | |||
+| (e 5.219) | |||
+| (e 5.220) |
 Hence
-| (e 5.207) |
-It follows from (e 5.205) and (e 5.206), and by the assumption on and that
-| (e 5.208) | |||
-| (e 5.209) |
+| (e 5.221) |
+It follows from (e 5.219) and (e 5.220), and by the assumption on and that
+| (e 5.222) | |||
+| (e 5.223) |
 Note Note also that has property (IR). By the choice of and applying Theorem 4.4 of [8], there is a normal element such that
-| (e 5.210) |
-Put Then is normal and, by (e 5.207) and (e 5.204),
-| (e 5.211) |
+| (e 5.224) |
+Put Then is normal and, by (e 5.221) and (e 5.218),
+| (e 5.225) |
 Put and Then, since is normal, Moreover
-| (e 5.212) |
+| (e 5.226) |
 The Proof of Theorem 1.3.
@@
 If is unital, then Therefore the conclusion follows from Theorem 4.4 of [8].
-Otherwise, by the assumption, has continuous scale (see of [27]).
+Otherwise, by the assumption, is simple.
+It follows that has continuous scale (see of [27]).
 Hence, by Corollary 3.3 of [27], is a simple purely infinite simple -algebra.
@@
 For there exists satisfying the following: Suppose that is an infinite dimensional separable Hilbert space and with () such that
-| (e 5.213) |
+| (e 5.227) |
 for all Then there are such that
-| (e 5.214) |
+| (e 5.228) |
 Proof.
-The next proposition provides a short way to see the second condition in (e 5.185) is necessary. We also provide a direct proof.
+The next proposition provides a short way to see the second condition in (e 5.199) is necessary. We also provide a direct proof.
 Proposition 5.5.
 Let There is satisfying the following: Suppose that and there are such that
-| (e 5.216) |
+| (e 5.230) |
 Then, for any
-| (e 5.217) |
-Note that (e 5.217) also implies that is invertible.
+| (e 5.231) |
+Note that (e 5.231) also implies that is invertible.
 Proof.
 Let be given by Proposition 2.15 for (in place of ). Choose Suppose that and are in the statement of the theorem, Put and Then
-| (e 5.218) |
+| (e 5.232) |
 Since is normal, is an essential -near-spectrum of By Proposition 2.15,
-| (e 5.219) |
+| (e 5.233) |
 where By (e 2.16), we may further have
-| (e 5.220) |
+| (e 5.234) |
 Since is normal, Note, for any
-| (e 5.221) |
-Now if then, by (e 5.220), Since is normal,
-| (e 5.222) |
-Hence, by (e 5.221),
-| (e 5.223) |
+| (e 5.235) |
+Now if then, by (e 5.234), Since is normal,
+| (e 5.236) |
+Hence, by (e 5.235),
+| (e 5.237) |
 This implies that is invertible and
@@
 Remark 5.6.
-There is a considerable overlap between the results in this section and that of Theorem 1.1 of [17] even though they have different proofs.
-However, Theorem 1.1 of [17] provides a quantitative estimate of the distance from to the set of normal elements with finite spectrum (perhaps the quantity is not ready to be computed) while Theorem 5.3 is a conceptional result.
-In the special case of (i) of Theorem 1.3, (defined in [17]) is equivalent to the condition that for all
-On the other hand, computationally speaking, -synthetic-spectrum could be tested.
-Note that a -synthetic-spectrum consists of finitely many balls and there are only finitely many bounded connected components in its complement.
-This allows one to avoid the situation of infinitely many “holes” in the spectrum.
-Note that the index in each of these “holes” remains constant.
-So one only needs to test one point in each of these “holes”.
-One also notes that the second condition in (e 5.185) does not imply
-It should also be noted that, when or a normal element ( is Theorem 5.3) may not be approximated by normal elements with finite spectrum even under the assumption that for all
-In Theorem 1.1 of [17], seems depending on a given operator The constant in Theorem 5.3 does not depend on nor even -algebras (as long as satisfies the conditions in Theorem 5.3).
-Finally, let us note that when a unital -algebra has (IR), then, by Theorem 4.4 of [8], a pair of almost commuting self-adjoint elements and is close to a pair of commuting ones even though may be large (see also the example below).
+Let us compare the results in this section and Theorem 1.1 of [17]:
+While there is significant overlap between the results in this section and Theorem 1.1 of [17], the proofs are quite different.
+Theorem 1.1 of [17] provides a quantitative estimate for the distance from to the set of normal elements with finite spectrum (though the explicit computation of may not be straightforward).
+In contrast, Theorem 5.3 here is a more conceptual result.
+In the special case of Theorem 1.3 (i), the condition (defined in [17]) is equivalent to for all
+Key Differences and Limitations:
+However, when or a normal element (as in Theorem 5.3) may not be approximable by normal elements with finite spectrum—even if for all In such cases, (from Theorem 1.1 of [17]) can be large.
+For instance, if is a unitary with in then rendering Theorem 1.1 of [17] inapplicable (see Example 5.7 for details).
+Moreover, if a unital -algebra has property (IR), then by Theorem 4.4 of [8], any pair of almost commuting self-adjoint elements and is close to a commuting pair – regardless of the size of (see also Example 5.7 below).
+Computational Aspects of -Synthetic Spectrum:
+From a computational standpoint, the -synthetic spectrum is tractable: it consists of finitely many balls, and its complement has only finitely many bounded connected components.
+Since the index remains constant in each such component, it suffices to test one point per bounded components.
+This avoids the complication of infinitely many “holes” in the spectrum
+Further points:
+The second condition in (e 5.199) does not imply
 Example 5.7.
@@
 Consider any unital -algebra which is given by the following essential extension:
+Let be the quotient map.
 Note that is purely infinite, has property (IR).
@@
 However, from the six-term exact sequence in -theory, one computes that Denote by the set of normal elements of and the set of those elements in which has finite spectrum, respectively.
-Then the closure Let be a non-zero projection.
-Since is either purely infinite or stable rank one, Thus there is a unitary with Let Then is a normal element with Hence for any Since and one sees easily that In other words, Theorem 1.1 in [17] could not be applied to determine when and are almost commuting (see also Theorem in [23]).
+Then the closure
+Let be a non-zero projection. Since is either purely infinite or stable rank one, Thus there is a unitary with Let Then is a normal element with Hence for any (so Theorem 1.3 (iii) applies).
+Then Choose Put Suppose that with Then, for any
+| (e 5.238) |
+It follows that In other words,
+Let be the disk in with center at and radius Then Choose such that if if if Define such that (). Note that and Then, there exists such that, for any pair of normal elements with and with then
+| (e 5.239) |
+Put
+Claim: but
+In fact, suppose otherwise there is such that
+| (e 5.240) |
+Then, since and
+| (e 5.241) |
+Hence and Therefore is a normal partial isometry. Moreover,
+| (e 5.242) |
+There exists a unitray such that
+| (e 5.243) |
+Also
+Note that is a unitary in with finite spectrum, hence This contradict the fact that in which proves the claim.
 6 Epilogue
-In this last section, we will address the relation between the condition of vanishing Fredholm index and that of closeness of the (synthetic)-spectrum and essential (synthetic)-spectrum (see Corollary 6.2 and 6.3).
-We will also give an example (Proposition 6.4) which shows the complexity behind the failure of a higher dimensional version of Theorem 1.3.
-A final remark on the Mumford original question related to quantum theory and measurements, and their relation with the second condition in Theorem 1.1 is also included.
+In this final section, we examine the connection between the vanishing Fredholm index condition and the closeness of the (synthetic) spectrum to the essential (synthetic) spectrum (Corollaries 6.2 and 6.3).
+We also present an example (Proposition 6.4) where a higher-dimensional analogue of Theorem 1.3 fails– highlighting the obstructions and subtleties behind this failure.
+Finally, we conclude with a remark on Mumford’s original question regarding quantum theory and measurements, relating it to the second condition in Theorem 1.1.
 Let us first state the following corollary of Theorem 4.6.
@@
 Let and There exists satisfying the following: Suppose that with () such that
-| (e 6.224) |
+| (e 6.244) |
 and Then there are such that
-| (e 6.225) |
+| (e 6.245) |
 Proof.
-The next corollary shows that, when the second condition in (e 4.156) is stronger than that of vanishing index in Theorem 1.3.
+The next corollary shows that, when the second condition in (e 4.170) is stronger than that of vanishing index in Theorem 1.3.
 Corollary 6.2.
 For any there exists satisfying the following: Let and be in all cases (i), (ii) and (iii) of Theorem 1.3. Suppose that
-| (e 6.226) |
+| (e 6.246) |
 Then, for any
-| (e 6.227) |
+| (e 6.247) |
 Proof.
 Let be given by Proposition 2.16 for
-We choose such that the conclusion of Proposition 2.15 holds for Then we choose such that the conclusion of Proposition 2.11 holds for Choose Suppose that (e 6.226) holds.
+We choose such that the conclusion of Proposition 2.15 holds for (in place of ).
+Then we choose such that the conclusion of Proposition 2.11 holds for Choose Suppose that (e 6.246) holds.
 Put Thus (applying Proposition 2.11) and are all non-empty sets.
 By Proposition 2.15,
-| (e 6.228) |
+| (e 6.248) |
 By applying Proposition 2.16, we have
-| (e 6.229) |
+| (e 6.249) |
 Applying (e 2.16), we obtain
-| (e 6.230) |
+| (e 6.250) |
 Since
-| (e 6.231) |
+| (e 6.251) |
 one has (see (e 2.16))
-| (e 6.232) |
-So, if By [5], in case (i) and (ii), since is stable, For case (iii), follows from Lemma 3.3 of [19] (note that, proved in [18], if has real rank zero, then ). Therefore
-| (e 6.233) |
+| (e 6.252) |
+So, if then By [5], in case (i) and (ii), since is stable, For case (iii), follows from Lemma 3.3 of [19] (note that, proved in [18], if has real rank zero, then ). Therefore
+| (e 6.253) |
 Hence
-| (e 6.234) |
-In general, we have the following statement.
-This follows immediately from Theorem 4.6.
-With the same spirit as that of the proof of Corollary 6.2, one could provide a direct proof without mentioning Theorem 4.6.
-However, since the point has been made in Corollary 6.2, we refrain ourselves from placing further -theory burden on the reader who might not be familiar with -theory.
+| (e 6.254) |
+In general, we have the following statement, which follows immediately from Theorem 4.6.
+Following the same spirit as the proof of Corollary 6.2, one could alternatively give a direct proof without invoking Theorem 4.6.
+However, since this approach is already demonstrated in Corollary 6.2, we avoid introducing additional -theoretic technicalities – particularly to spare readers less familiar with -theory.
 Corollary 6.3.
 Let For any there exists satisfying the following: Suppose that with () such that
-| (e 6.235) |
+| (e 6.255) |
 where and Then there exists a -brick combination and a unital monomorphism such that
-| (e 6.236) | |||
-| (e 6.237) |
+| (e 6.256) | |||
+| (e 6.257) |
 Proof.
@@
 Since for any compact metric space we conclude that Consequently where ∎
-We now present an example which shows that an -dimensional version of part (i) of Theorem 1.3 is false ().
-Note that, in the next proposition, the triple has a single point in the essential -spectrum.
-Hence the -theory obstruction vanishes in
+We now present a counterexample demonstrating that the -dimensional analogue of Theorem 1.3 (i) is false when Note that, in the next proposition, the triple has only a single point in the essential -spectrum.
+Hence the -theory obstruction vanishes in This example also underscores the significance of Theorems 1.1 and 1.2.
 Proposition 6.4.
 There exists and a sequence of triples of self-adjoint operators such that (for all ),
+where infimum is taken from all possible commuting triples of selfadjoint operators and all
 Proof.
 Let be the three dimensional unit ball in By Theorem 4.1 of [9], there exists a sequence of contractive linear maps and which satisfy the following:
-| (e 6.238) | |||
-| (e 6.239) |
+| (e 6.258) | |||||
+| (e 6.259) |
 where the infimum is taken among all homomorphisms
@@
 Let Then there exists a sequence of projections such that Put Note that
-| (e 6.240) |
+| (e 6.260) |
 Hence
-| (e 6.241) |
+| (e 6.261) |
 There are such that, for all sufficiently large
-| (e 6.242) |
+| (e 6.262) |
 Therefore, by replacing with we may assume that maps into and each is unital.
-Choose Identify as a finite rank () projection. Put Define by
-| (e 6.243) |
-Choose
-| (e 6.244) |
+Choose Identify with a -subalgebra of such that becomes a projection of rank In particular, as a finite rank () projection. Put Define by
+| (e 6.263) |
+Choose (with )
+| (e 6.264) |
 So and Moreover
-| (e 6.245) |
+| (e 6.265) |
 Furthermore,
-| (e 6.246) |
+| (e 6.266) |
+where is the quotient map.
 Let and be the quotient map. Define by for
 If there is a sequence of commuting triples of self-adjoint operators such that
-| (e 6.247) |
+| (e 6.267) |
 then one obtains a sequence of unital homomorphisms such that
-| (e 6.248) |
+| (e 6.268) |
 Hence
-| (e 6.249) |
+| (e 6.269) |
 Then where is defined by for
 Choose such that if and if Note that and
-| (e 6.250) |
+| (e 6.270) |
 Put and Then
-| (e 6.251) |
+| (e 6.271) |
 Recall that
-| (e 6.252) |
+| (e 6.272) |
 Put and It follows that
-| (e 6.253) |
+| (e 6.273) |
 Let be the (commutative) -subalgebra of generated by Then there is an isomorphism for some compact subset such that and Define by for Since generates we have
-| (e 6.254) |
+| (e 6.274) |
 Let Then,
-| (e 6.255) |
+| (e 6.275) |
 Thus, for all large there exist unitaries such that
-| (e 6.256) |
+| (e 6.276) |
 Define by for Then
-| (e 6.257) |
-This contradicts (e 6.239). Therefore there is no homomorphisms such that (e 6.249). In other words, there is such that
-| (e 6.258) |
-where the infimum is taken among all homomorphisms Combining this with (e 6.246), the proposition follows. ∎
+| (e 6.277) |
+This contradicts (e 6.259). Therefore there is no homomorphisms such that (e 6.269) holds. In other words, there is such that
+| (e 6.278) |
+where the infimum is taken among all homomorphisms Combining this with (e 6.266), the proposition follows. ∎
 Remark 6.5.
 Remark 6.6.
-This remark was added to the original version of the paper after the author fortunately received a copy of Chapter 14 of a very exciting new book of David Mumford [34] graciously sent by the author.
-Some terminologies are taken from there.
-In quantum theory, one may view macroscopic observables as some self-adjoint operators on a Hilbert space which we assume to be infinite dimensional and separable.
-Let be the unital -subalgebra generated by and (let us assume for convenience here).
-Suppose that is an orthonormal basis of and suppose that, for any there is such that for all In this case these observables collapse uniformly and any measurements outside (orthogonally) a finite dimensional subspace vanish.
-In other words, outsiders would see nothing.
-Let us assume that none of these observables collapse this way (see also Remark 6.5).
-In other words, we assume that none of ’s are compact.
-Given a state which is defined by a unit vector the expected value when measuring an observable given by an operator is (see p.178 of [34]).
-“Approximately Macroscopically Unique” states introduced by Mumford (see Chapter 14 of [34]), or AMU states, are “near eigenvectors”.
-Set, for a small
-where As pointed out in [34], for AMU to be non-empty, it is natural to assume that the commutators have small bound.
-If the question in the introduction (see (e 1.1) and (e 1.2)) had an affirmative answer, then the quantum system could be approximated by a classical one.
-Consequently, AMU does exist.
-Observed by Mumford, the condition that the commutators are small is not quite enough (Proposition 6.4 further shows the complexity of the problem) in general.
-In an infinite dimensional separable Hilbert space a self-adjoint operator such as may not have any eigenvalues (but have approximate eigenvalues).
-However, if one allows non-normal states of or that of then does have eigenstates and eigenvalues in the atomic representation of or (see, for example, Proposition 4.3.10 of [35]).
-In particular, one should consider those states of which vanish on the ideal of compact operators.
-If a point in the unit cube of is expected to be a “joint eigenvalue”, then one may use the test function defined in 2.7 to test it.
+This remark was added to the original version of the paper after we had the opportunity to study Chapter 14 of David Mumford’s insightful new book [34], kindly provided by the author.
+Some terminologies and motivation are drawn from this work.
+In quantum theory, macroscopic observables may be modeled as self-adjoint operators on an infinite dimensional separable Hilbert space Let be the unital -subalgebra generated by and (assuming for convenience).
+Suppose that is an orthonormal basis of and suppose that, for any there is such that for all In this case these observables collapse uniformly and any measurements outside (orthogonal to) a finite dimensional subspace vanish, meaning an external observer would detect nothing.
+We exculde this case by assuming none of ’s are compact (see also Remark 6.5).
+Given a state defined by a unit vector the expected value of observable is (see p.178 of [34]). Mumford introduces Approximately Macroscopically Unique (AMU) states as “near eigenvectors”: for a small
+where (see Chapter 14 of [34]).
+Mumford asked whether AMU exist.
+As noted in [34], for AMU to be non-empty, it is natural to require the commutators to be small.
+If the question posed in the introduction (see (e 1.1) and (e 1.2)) had an affirmative answer, then the quantum system could be approximated by a classical one, ensuring the existence of AMU states.
+However, as observed by Mumford, small commutators alone are insufficient in general.
+(Proposition 6.4 further shows the complexity of the problem).
+In an infinite dimensional separable Hilbert space a self-adjoint operator such as may lack eigenvalues but admit approximate eigenvalues.
+However, if one allows non-normal states of or that of (where is the ideal of compact operators), then does have eigenstates and eigenvalues in the atomic representation of or (see, for example, Proposition 4.3.10 of [35]).
+In particular, one should consider those states of which vanish on
+For a point in the unit cube of to be a joint approximate eigenvalue, we may test it using the function defined in 2.7.
 In order to have (for small ), one only needs one vector state (with a unit vector ) such that
-| (e 6.259) |
+| (e 6.279) |
 One may expect this can be repeatedly measured.
-Suppose that are unit vectors and is a finite dimensional subspace spanned by Suppose that there is another unit vector such that (e 6.259) also holds for (in place of ), and this persists.
-Then we obtain a sequence of () such that (e 6.259) holds for each (in place of ).
-Note that unit vector states given by has limit points which give states of vanishing on and Note that there are still unknown (approximate) eigenvalues outside any finite dimensional subspaces of
+Suppose that are unit vectors and is a finite dimensional subspace spanned by Suppose that there is another unit vector such that (e 6.279) also holds for (in place of ), and this persists.
+Then we obtain a sequence of () such that (e 6.279) holds for each (in place of ).
+Note that unit vector states given by has limit points which give states of vanishing on and
 Let be the quotient map and be a state of the corona algebra (or that of ).
-Then the quantum state might better reflect this situation.
-Indeed, one may expect that, at least for one such quantum state one still has
-| (e 6.260) |
-Measurements given by states of the form may be called outside measurements.
-By the Gelfand-Naimark construction, this quantum state is also given by a vector state (the Hilbert space from the atomic representation of or ).
-In other words,
-| (e 6.261) |
-If whenever (e 6.259) holds, inequality (e 6.260) holds for some quantum states then the second condition in (e 1.3) always satisfies (in fact, in this case ).
-Let us, for a moment, call a unit vector state for which (e 6.259) holds a synthetic-AMU state and a synthetic-joint-expected value.
-The state for which (e 6.260) holds may be called an essential synthetic-AMU and (for which (e 6.260) holds) an essential synthetic-joint-expected value.
-Note that it is easy to see an AMU state (with sufficiently small ) is an ()-synthetic-AMU state.
-Moreover, by Proposition 2.11, the set of synthetic-AMU states is non-empty, whenever is small (independent of ).
-The second condition in Theorem 1.1 then may be stated as follows:
-(C): all synthetic-joint-expected values are in a sufficiently small neighborhood of the set of essential synthetic-joint-expected values.
-This seems to suggest that we could put a condition on the macroscopic system: any local measurement should not be too far off from an outside (or a long-term measurements along an orthonormal basis of the Hilbert space) measurement.
-Under this condition, when commutators tend to zero, by Theorem 1.1, the quantum system may be approximated by a classical system where observables commute.
+Then the composition provides a quantum state which may be better suitable to this setting.
+Indeed, one may expect at least for one such quantum state satisfying
+| (e 6.280) |
+Such measurements, given by states of the form may be called outside measurements.
+Via the Gelfand-Naimark construction, this state corresponds to a vector state in the atomic representation space yielding
+| (e 6.281) |
+If inequality (e 6.280) holds whenever (e 6.279) does, then the second condition in (e 1.3) always satisfies (in fact, in this case ).
+Let us temporerly propose the following terminology:
+A unit vector satisfying (e 6.279) is a synthetic-AMU state, with a synthetic-joint-expected value.
+A state of the form satisfying (e 6.280) is an essential synthetic-AMU state, with an essential synthetic-joint-expected value.
+Note that any AMU state (with sufficiently small ) is an ()-synthetic-AMU state.
+Moreover, by Proposition 2.11, the set of synthetic-AMU states is non-empty, when is small (independent of ).
+The second condition in Theorem 1.1 then may be reformulated as:
+(C): all synthetic-joint-expected values lie in a sufficiently small neighborhood of the set of essential synthetic-joint-expected values.
+This seems to suggest a physical constraint on macroscopic systems: local measurement should not deviate significantly from outside measurements (or a long-term measurements along an orthonormal basis of the Hilbert space).
+Under this constraint, when commutators tend to zero, by Theorem 1.1, the quantum system becomes approximable by a classical commutative one.